### v1

# Dynamic Mixture of Progressive Parameter-Efficient Expert Library for Lifelong Robot Learning

Abstract
A generalist agent must continuously learn and adapt throughout its lifetime, achieving efficient forward transfer while minimizing catastrophic forgetting. Previous work within the dominant pretrain-then-finetune paradigm has explored parameter-efficient fine-tuning for single-task adaptation, effectively steering a frozen pretrained model with a small number of parameters. However, in the context of lifelong learning, these methods rely on the impractical assumption of a test-time task identifier and restrict knowledge sharing among isolated adapters. To address these limitations, we propose Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL) for lifelong robot learning. DMPEL progressively learn a low-rank expert library and employs a lightweight router to dynamically combine experts into an end-to-end policy, facilitating flexible behavior during lifelong adaptation. Moreover, by leveraging the modular structure of the fine-tuned parameters, we introduce coefficient replay to guide the router in accurately retrieving frozen experts for previously encountered tasks, thereby mitigating catastrophic forgetting. This method is significantly more storage- and computationally-efficient than applying demonstration replay to the entire policy. Extensive experiments on the lifelong manipulation benchmark LIBERO demonstrate that our framework outperforms state-of-the-art lifelong learning methods in success rates across continual adaptation, while utilizing minimal trainable parameters and storage.
1 Introduction
A generalist agent should have the capability to learn and adapt continuously throughout its lifetime, usually termed as lifelong learning [1, 2, 3, 4, 5, 6]. The longstanding challenges are enabling forward transfer (i.e., leveraging knowledge from previous tasks to quickly adapt to new ones) and avoiding catastrophic forgetting (i.e., retaining previously acquired knowledge when learning new tasks) under limited computational and memory capacity.
Due to the notorious sample inefficiency of the traditional tabula rasa approach in robotics, researchers have recently explored the pretrain-then-finetune paradigm [7, 8, 9, 10, 11, 12] that originates from the vision and language domains [13, 14, 15, 16, 17, 18, 19, 20]. Specifically, this paradigm first pretrain a policy, often referred to as a vision-language-action (VLA) model, on large-scale datasets, and then fine-tune it for various downstream tasks. However, in the context of lifelong learning, naively applying full fine-tuning to sequentially arriving robotic tasks can result in severe forgetting and suboptimal performance, failing to meet our requirements. As depicted in Figure 1, prior work in lifelong robot learning can be broadly classified into three distinct approaches. Replay methods [8, 21, 22, 23, 24, 25] retain previous data (e.g., vision-language web data, robot pretraining data, and data from seen tasks) and mix it with in-domain data during training on new tasks. These methods leads to generalizable policies, but also necessitates huge storage space and computational overhead for retaining old samples. Regularization methods [26, 27, 28] balance between new and old tasks by restricting the update of parameters. For example, EWC [27] require an entire copy of frozen model from the previous task and entail substantial computation in estimating the importance of parameters. In addition, they usually struggle in long task sequence due to plasticity-stability dilemma.
In contrast, architectural methods [29, 30, 31] address catastrophic forgetting by explicitly learning task-specific parameters. The widely adopted parameter-efficient fine-tuning (PEFT) techniques [32, 33], such as adapters [34, 35], low-rank adaptation (LoRA) [36], and prompt tuning [37, 38, 39], have proven highly effective for steering frozen foundation models by integrating small, learnable modules in single-task adaptation. In the lifelong learning setting, a practical solution, as exemplified by TAIL [40], involves initializing and training task-specific adapters for each task, subsequently freezing them, and retrieving the appropriate adapter based on the oracle task index. Such method allows efficient adaptation with only a small amount of task-specific parameters, but also exhibits certain limitations. First, assuming test-time task identity is known and retrieving adapter in a hard-coded way could be impractical in real scenarios. Second, failing to effectively share knowledge across isolated adapters result in poor forward transfer. Some prior arts construct a shared adapter pool and enable automatic timestep-level retrieval by, for example, performing query-key matching to select the one with highest similarity from a pool of adapters [41], or using multifaceted prototypes to represent the cluster center in the state space for each adapter [42], but usually suffer from inaccurate retrieval and suboptimal performance.
Figure 1: Existing lifelong learning methods: (a) Replay methods; (b) Regularization methods; (c) Architectural methods; (d) Performance on LIBERO-Object. TAIL with LoRA utilizes significantly fewer resources than ER/EWC with FFT and exhibits no forgetting when provided with a task identifier, but demonstrates lower forward transfer. In contrast, DMPEL leverages a low-rank expert library and coefficient replay (CR$=$5%) to achieve better forward transfer and near-zero forgetting.
In this paper, we introduce Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL) to overcome these challenges in lifelong robot learning. DMPEL incrementally builds a library of low-rank experts and uses a lightweight router to dynamically assemble these experts into a unified policy, enabling the agent to adapt flexibly throughout continual adaptation. By exploiting the modularity of fine-tuned parameters, DMPEL incorporates coefficient replay at the router, enabling the policy to accurately integrate experts for all previously seen tasks and thereby effectively preventing catastrophic forgetting. This approach is significantly more efficient in terms of storage and computation than replaying expert demonstrations for the entire policy. Through extensive evaluations on the LIBERO lifelong manipulation benchmark [43], we show that our method achieves higher forward transfer and lower catastrophic forgetting than existing baselines while requiring minimal trainable parameters and storage. Code is publicly released for reproducibility.
2 Related Work
Foundation Models for Robotics.
The traditional tabula rasa paradigm in robotics is widely criticized for its sample inefficiency, prompting researchers to explore the use of foundation models to improve downstream task transfer [44, 45]. One prominent direction involves transferring general visual representations: some studies directly employ foundation models from other domains, such as CLIP [15], SAM [18], and DINOv2 [14], as vision encoders [46, 47], while others pretrain visual representations using self-supervised learning [16, 17] on egocentric video datasets like R3M [48], MVP [49], and VC-1 [50]. Another emerging line of research focuses on pretraining large-scale policies on diverse robotic tasks to capture generalizable behavior priors, for example, RT-1 [7], RT-2 [8], Gato [9], Octo [10], OpenVLA [11], and $\pi_{0}$ [12]. However, when adapting these models to downstream tasks, most approaches either rely on the frozen pretrained model (i.e., zero-shot transfer) or require costly full fine-tuning. Some prior work has explored parameter-efficient fine-tuning (PEFT) by integrating task-specific modules into the pretrained vision encoder [51, 52, 53] or temporal transformer [54, 55, 56]. Despite these advancements, the majority of these methods focus on single-stage adaptation, whereas our approach targets lifelong adaptation to a sequence of new tasks.
Lifelong Learning with PEFT.
Compared to full fine-tuning, PEFT optimizes only a small amount of inserted parameters and keeps the rest untouched, greatly reducing the storage cost of tuned parameters and mitigating the risk of overfitting in low-data regimes [32]. Notably, PEFT techniques share conceptual similarities with architectural methods in lifelong learning. Most studies on lifelong learning with PEFT predominantly utilize prompt-based tuning [37, 39] to adapt frozen pretrained model for tasks such as image classification (e.g., L2P [57], DualPrompt [58], CODA-Prompt [59], and DAP [60]) and text classification (e.g., EPI [61], and Progressive Prompts [62]). Current research primarily investigates strategies for dynamically selecting or generating instance-specific prompts [63]. Recognizing the representational limitations of prompt tuning, subsequent studies have extended to more powerful PEFT techniques, including LAE [64], HiDe-PET [65]. Despite the demonstrated success of PEFT-based lifelong learning in classification tasks, its application remains largely unexplored in robotics. A few pioneering works employ LoRA [36] to store task-specific knowledge, yet diverge in their retrieval mechanisms: TAIL [40] relies on oracle task identifier, L2M [41] implements query-key matching between the current context (the state embedding from the frozen pretrained model) and learnable keys, while IsCiL [42] establishes multifaceted prototypes through K-means clustering in the state space, with input states retrieving the closest LoRA parameters based on proximity. Our method incrementally constructs a parameter-efficient expert library and learns a lightweight router to dynamically retrieve and combine these experts.
Mixture of Experts (MoE) and Model Fusion.
The Mixture of Experts (MoE) approach, which involves training multiple specialized models for specific subtasks, has gained significant attention [66, 67, 68]. A key method for combining these experts is parameter ensemble, supported by the linear mode connectivity phenomenon, which shows models converging to a single low-loss basin [69]. Task Arithmetic [70] introduces task vectors, representing the difference between fine-tuned and pretrained models, enhancing multi-task model performance when merged. Follow-up work [71, 72] improves this by learning input-conditioned layer-wise fusion coefficients, increasing flexibility and performance. Model fusion techniques also apply to PEFT, where previously learned adapters are reused for efficient task adaptation [73, 74, 75, 76], including LoRAHub [77] and AdapterFusion [78]. In image generation, merging separate LoRAs facilitates joint style and content generation [79, 80, 81]. In robotics, models like MELA [82], MoE-Loco [83], MORE [84], and SDP [85] create versatile, adaptive behaviors for multiple tasks by fusing expert sets. Our method aims to enhance forward transfer in lifelong robot learning by flexibly fusing previously learned low-rank experts for each sub-module in the policy. Besides, we exploit the modularized design by replaying coefficient on the router to mitigate catastrophic forgetting.
3 Problem Formulation
3.1 Lifelong Robot Learning
In this paper, we aim to solve language-conditioned vision-based robot manipulation task, which can be formulated as a finite-horizon Markov Decision Process (MDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{H},d_{0},l)$, where $\mathcal{S}$ and $\mathcal{A}$ is the state space and action space of robot, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition dynamics, $\mathcal{H}$ is the episode length, and each task $\mathcal{T}\triangleq(d_{0},l)$ is defined by the initial state distribution $d_{0}$ and the goal predicate in the form of language instruction $l:\mathcal{S}\rightarrow\{0,1\}$. The ultimate objective of robot learning is to search for a policy $\pi$ that maximizes the expected success rate in reaching the goal: $\max_{\pi}J(\pi)=\mathbb{E}_{s_{t},a_{t}\sim\pi,d_{0}}\left[\sum_{t=1}^{\mathcal{H}}l(s_{t})\right]$.
Due to the significant challenge of sparse reward in robot manipulation, we consider a more practical lifelong imitation learning setting [43, 40], where a robot sequentially learns over a stream of tasks $\{\mathcal{T}^{1},\cdots,\mathcal{T}^{K}\}$ given an expert demonstration dataset $\mathcal{D}^{k}=\{\tau_{n}^{k}\}_{n=1}^{N}$ for each task $\mathcal{T}^{k}=(d_{0}^{k},l^{k})$, $1\leq k\leq K$. Each expert trajectory $\tau_{n}^{k}$ can be represented as $(o_{0},a_{0},\cdots,o_{\mathcal{H}})$, in which observation $o_{t}$ at each step $t$ includes RGB images and proprioceptive states. Following common pratice in partially observable MDPs, we approximate the current state $s_{t}$ by stacking prior observations, i.e., $s_{t}=o_{\leq t}=(o_{0},\cdots,o_{t})$. We perform behavior cloning [86] to learn a stochastic policy $\pi_{\theta}$ by minimizing the negative log-likelihood loss $L$:
$\min_{\theta}L(\theta)=-\mathbb{E}_{\tau_{n}^{k}\sim\mathcal{D}^{k}}\left[\sum_{t=0}^{\mathcal{H}-1}\log\pi_{\theta}\left(a_{t}|o_{\leq t},l^{k}\right)\right].$ (1)
3.2 Evaluation Metrics
We use three metrics to evaluate the performance of lifelong robot learning [43, 87]: forward transfer (FWT), negative backward transfer (NBT), area under the success rate curve (AUC). Formally, we evaluate the policy on current task $\mathcal{T}^{k}$ when reaching a checkpoint $c\in\mathcal{C}$ (i.e., every multiple epochs of training) and we denote the success rate during evaluation as $s_{k,k,c}$. When the training on current task is completed, we find the earliest checkpoint $c_{k}^{*}$ that achieves the best performance and mark it as the final success rate $s_{k,k}$, i.e., $s_{k,k}=s_{k,k,c_{k}^{*}}=\max_{c\in\mathcal{C}}s_{k,k,c}$. We only keep the best checkpoint for future training and evaluation as if the agent stops learning after $c_{k}^{*}$, i.e., for all $c\geq c_{k}^{*}$, $s_{k,k,c}=s_{k,k}$. This checkpoint $c_{k}^{*}$ will be further evaluated on previous tasks $\mathcal{T}^{j},j\leq k$, where the success rate is denoted as $s_{k,j}$. Then we can define the three metrics as:
FWT $=\frac{1}{K}\sum_{k=1}^{K}\left[\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}s_{k,k,c}\right],$ (2)
NBT $=\frac{1}{K}\sum_{k=1}^{K}\left[\frac{1}{K\!-\!k}\sum_{l=k+1}^{K}\left(s_{k,k}-s_{l,k}\right)\right],$ (3)
AUC $=\frac{1}{K}\sum_{k=1}^{K}\left[\frac{1}{K\!-\!k\!+\!1}\Big{(}\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}s_{k,k,c}+\!\!\sum_{l=k+1}^{K}\!\!s_{l,k}\Big{)}\right].$ (4)
Intuitively, higher FWT indicates faster adaptation to new tasks, lower NBT indicates less catastrophic forgetting on old tasks, and higher AUC demonstrates an overall better performance across tasks and a good balance between FWT and NBT.
4 Proposed Method
4.1 Base Policy Architecture
We use an autoregressive policy [11, 88] similar to the one used in the LIBERO benchmark [40, 43] and focus on developing better lifelong learning algorithms. The policy consists of vision, text, and state encoders, an input modality fusion module, a temporal transformer, and an action head.
Vision, text, and state encoders.
The policy input is the historical observations $(o_{t-T},\cdots,o_{t})$ and the language instruction $l=\{l_{i}\}_{i=1}^{L}$, where $T$ is the context length and $L$ is the sentence length. The observation at each step $o_{t}=(I^{1}_{t},\cdots,I^{N_{c}}_{t},p_{t})$ includes RGB images $I^{n_{c}}_{t},1\leq n_{c}\leq N_{c}$ from $N_{c}$ different cameras and low-dimensional proprioceptive states $p_{t}$. In experiments, we use images from two cameras, the agent-view image and the eye-in-hand image. We employ the pretrained CLIP ViT-B/16 model [15] to serve as the vision encoder $\mathcal{E}_{I}$ and the text encoder $\mathcal{E}_{L}$. For the joint states and gripper states, we learn separate linear projection layers $\mathcal{E}_{P}$ to project the low-dimensional states to proprioceptive embeddings. In summary, at each timestep, we obtain image embeddings $f^{v}_{t}$, text embeddings $f^{l}_{t}$, and proprioceptive embeddings $f^{s}_{t}$.
Modality fusion.
We utilize a feature-wise linear modulation (FiLM) [89] to fuse language embeddings with the image and state embeddings. Generally, given the original feature $x$, the conditional input $z$, and the fusion network $f_{\text{FiLM}}$, we can obtain the modulated feature $x^{\prime}=\gamma\odot x+\beta$ where $\gamma,\beta=f_{\text{FiLM}}(z)$. In our case, the conditional input $z$ refers to the language embedding $f^{l}_{t}$, $\gamma$ and $\beta$ are scaling and shifting vectors having the same size as $x$, and $\odot$ refers to element-wise multiplication.
Temporal transformer and action head.
The causal temporal transformer backbone processes a sequence of multi-modal embeddings to produce a latent vector $g_{t}$ at each decision-making timestep. We employ a Gaussian Mixture Model (GMM)-based output head to compute the multi-modal distribution of manipulation actions [40, 43]. During training, we optimize the negative log-likelihood loss between the action distribution and the ground truth action. At inference time, the robot executes the policy by using the mean of the Gaussian distribution with the highest density for end effectors.
Figure 2: Overview of the proposed method DMPEL.
4.2 Low-rank Expert Library
Following the pretrain-then-finetune paradigm, we pretrain the policy on a large-scale robotic dataset and freeze it during the lifelong adaptation to unseen downstream tasks.
Low-Rank Adaptation (LoRA) [36].
A representative parameter-efficient fine-tuning (PEFT) method that introduces learnable low-rank matrices $\boldsymbol{A}\in\mathbb{R}^{d_{\text{in}}\times r}$, $\boldsymbol{B}\in\mathbb{R}^{r\times d_{\text{out}}}$ and integrate in parallel with the frozen weight matrix $\boldsymbol{W}_{0}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ through addition, i.e., $\boldsymbol{W}_{0}+\boldsymbol{A}\boldsymbol{B}$. By leveraging the low-rank decomposition $r\ll\min\{d_{\text{in}},d_{\text{out}}\}$, LoRA substantially reduces the number of trainable parameters.
Existing PEFT methods learn a task-specific low-rank adapter $\Delta\boldsymbol{W}_{k}=\boldsymbol{A}_{k}\boldsymbol{B}_{k}$ for each task $\mathcal{T}_{k}$ in lifelong robot learning [40, 41]. Such design requires task identifier during inference, lacks flexibility due to fixed capacity, and also hinders knowledge sharing across tasks. Prior arts in vision and language domain, e.g., LoRAHub [77] and SD-LORA [90], use a task-wise learnable weight vector to combine existing low-rank experts for better performance on the new task. A key insight is that these low-rank experts captures complementary knowledge from a variety of specific tasks and can be leveraged to solve a new task. However, lifelong learning in robotic domain could be even more challenging. First, compared to image or text classification [1, 3, 5] that requires only single forward inference, robotic manipulation is a sequential decision-making process with long horizon. Besides, robot learning requires a mixed types of knowledge, for example, visual concepts, textual task goals, different spatial relationships, and successful adaptation to a new task requires a combination of adaptation in each sub-modules.
Following this motivation, we propose to learn a progressive low-rank expert library on top of the pretrained policy and dynamically reuse low-rank experts from prior tasks $j\leq k-1$ to facilitate efficient forward transfer to new task $\mathcal{T}_{k}$. We equip each chosen pretrained linear layer =$\mathcal{F}=\{\boldsymbol{W}_{0},\boldsymbol{b}_{0}\}$ with a low-rank expert library =$\mathcal{L}=\{\boldsymbol{A}_{j},\boldsymbol{B}_{j},\boldsymbol{b}_{j}\}_{j=1}^{k}$ and use a global lightweight router $\mathcal{R}$ to dynamically integrate pretrained and task-specific knowledge based on the current context, providing a more flexible solution for adaptation.
Specifically, the router $\mathcal{R}:\mathbb{R}^{d_{r}}\rightarrow\mathbb{R}^{6\times k}$ is implemented as a multi-layer perceptron (MLP) to obtain the dynamic coefficient for each low-rank expert. The router input $\boldsymbol{r}_{t}=[f^{v}_{t-T:t},f^{l},p_{t-T:t}]\in\mathbb{R}^{d_{r}}$, where $\cdot_{t-T:t}$ indicates mean pooling over the $T$-step context window. This aggregated representation contains information streams from the visual embeddings, the task instruction embeddings, and also the low-dimensional proprioceptive states, and hence comprehensively encodes the current context. Then we can dynamically obtain the expert coefficient:
$\boldsymbol{c}_{t}=[\boldsymbol{c}^{v}_{t};\boldsymbol{c}^{l}_{t};\boldsymbol{c}^{s}_{t};\boldsymbol{c}^{f}_{t};\boldsymbol{c}^{t}_{t};\boldsymbol{c}^{h}_{t}]=\text{top-k}(\mathcal{R}(\boldsymbol{r}_{t}),\delta),\quad\boldsymbol{c}_{t}^{\cdot}=[c_{1,t}^{\cdot},c_{2,t}^{\cdot},\cdots,c_{k,t}^{\cdot}]\in\mathbb{R}^{k}.$ (5)
[…]
$\boldsymbol{y}=\tilde{\boldsymbol{W}}\boldsymbol{x}+\tilde{\boldsymbol{b}}=\underbrace{\boldsymbol{W}_{0}\boldsymbol{x}+\boldsymbol{b}_{0}}_{\text{pretrained}}+\underbrace{\sum_{j=1}^{k}c_{j,t}\boldsymbol{A}_{j}\sum_{j=1}^{k}c_{j,t}\boldsymbol{B}_{j}\boldsymbol{x}+\sum_{j=1}^{k}c_{j,t}\boldsymbol{b}_{j}}_{\text{fine-tuned}}.$ (6)
Upon the completion of task $\mathcal{T}_{k}$, we freeze all low-rank experts in the library as acquired knowledge and initialize new learnable expert for task $\mathcal{T}_{k+1}$. In order to reduce the interference between tasks, we employ the Gram-Schmidt orthogonalization method on $\boldsymbol{A}_{k+1}$ [59, 91], while other elements $\boldsymbol{B}_{k+1}$ and $\boldsymbol{b}_{k+1}$ are simply zero-initialized.
4.3 Expert Coefficient Replay
To mitigate catastrophic forgetting, we freeze the learned low-rank expert module after each task’s completion for future retrieval. However, since the router is continuously updated on sequentially arriving tasks, incorrect integration of the expert, i.e., the expert coefficients are inaccurate, may lead to severe policy distortion and suboptimal performance.
Inspired by replay methods [22, 24, 76], we propose an expert coefficient replay mechanism specifically designed for the lightweight router. The implementation involves two phases: During task finalization, we archive a subset of the router’s input-output pairs, i.e., the context embedding $\boldsymbol{r}$ and the coefficient vector $\boldsymbol{c}$, in a dedicated buffer $\mathcal{B}_{\text{CR}}$. When adapting to new tasks, the core idea is to enforce consistency by requiring the router to generate coefficients that closely match the stored historical coefficients given the same context. This is achieved through minimizing a mean squared error (MSE) objective:
$L_{\text{CR}}(\mathcal{R})=\mathbb{E}_{{(\boldsymbol{r}_{i},\boldsymbol{c}_{i,\text{old}})\sim\mathcal{B}_{\text{CR}}}}\frac{1}{2}\left(\mathcal{R}(\boldsymbol{r}_{i})-\boldsymbol{c}_{i,\text{old}}\right)^{2}.$ (7)
Compared with conventional expert demonstration replay that requires computation through the entire policy, our coefficient replay strategy demonstrates superior efficiency by focusing exclusively on the lightweight router and its inputs and outputs, yielding substantial reductions in both computational overhead and storage requirements.
5 Experiments
5.1 Experimental Setup
Benchmark.
We conduct evaluations in LIBERO [43], a lifelong learning benchmark on robot manipulation tasks. The robot in the benchmark is situated in a tabletop environment and equipped with a 6-DOF arm and a parallel gripper. We use four lifelong learning task suites: Goal, Spatial, Object, and Long. Each suite consists of 10 robotic manipulation tasks designed to investigate the transfer of knowledge related to task goals, spatial information, and various objects. Details can be found in Appendix A.3. The benchmark also includes LIBERO-90, a collection of 90 short-horizon diverse tasks that serves as a pretraining dataset for downstream transfer.
Baselines.
We consider a naive sequential fine-tuning baseline, three representative lifelong learning methods that involve tuning the entire policy, and several state-of-the-art LoRA-based methods:
  • Sequential Fine-tuning (SeqFT) [43]: Naively perform full fine-tuning (FFT) or use frozen pretrained feature (FPF), i.e., freeze the spatial encoders and fine-tune the rest, on sequentially arriving robot learning tasks.
  • Experience Replay (ER) [22], a replay method that maintains a memory buffer with samples from previous tasks. We store a subset of data after each task and uniformly sample a fixed number of replay data from the buffer along with new task data during training.
  • Elastic Weight Consolidation (EWC) [27], a regularization method that uses the Fisher Information Matrix (FIM) to quantify the importance of each parameter and adds a regularization term to the learning objective to constrain network updates.
  • PackNet [30], an architectural method that first trains and prunes the network to retain a fixed proportion of the most important parameters. The selected parameters are then fine-tuned and frozen to prevent forgetting.
  • LifelOng knowledge Transfer Using Skills (LOTUS) [92] involves two stages: extracting temporal segment features by leveraging frozen foundation models for skill discovery, and hierarchical imitation learning with the meta-controller and a growing skill library.
  • Task-specific Adapters for Imitation Learning (TAIL) [40] introduces separate LoRAs into the spatial encoder and temporal decoder for each new task. The fusion module, state encoder, and action head are also full fine-tuned and stored as task-specific adapters. TAIL requires an oracle task identifier to retrieve the corresponding adapters during inference. In order to align with other methods, we also provide the result of SeqFT (LoRA), which fine-tunes LoRA parameters sequentially.
  • Learning to Modulate (L2M) [41] maintains a learnable modulation pool with keys and associated modulators. The embedded history of state tokens serves as the query vector for query-key matching, which finally steers the pretrained policy’s behavior. We follow the original paper to use a frozen fusion module and action head, and insert LoRA pools in the CLIP encoders and the temporal transformer backbone.
  • Incremental Learning of Retrievable Skills for Continual Imitation Learning (IsCiL) [42] establishes multifaceted prototypes through K-means clustering in the state space, where proper LoRA-based skills is retrieved upon current input state. We define the task-specific skill the same as in TAIL [40], i.e., LoRA matrices for transformer, the fusion module, the state encoder, and the action head.
Implementation Details.
We utilize the open-source CLIP ViT/B-16 model [15] as vision and text encoders, while all other components of the policy is learned from robot demonstrations. We use a six-layer temporal transformer, using the same embedding size as the CLIP model $d=768$. The LoRA rank is set to 8 for CLIP encoders and 16 for others, the same as in [40]. We pretrain the base policy on the LIBERO-90 dataset and use it as the starting point of lifelong learning. During adaptation, we learn 10 epochs on each arriving task and evaluate every 2 epochs, using a batch size of 32 and AdamW optimizer with a learning rate of 1e-4. We perform expert coefficient replay for 10 epochs after each task. As for the expert router, we employ a scaled sigmoid output activation to restrict the coefficient in $[0,2]$ and use a top-3 strategy to select useful experts. All results are averaged over three random seeds. More details can be found in Appendix A.
Figure 3: Performance of the pretrained policy on LIBERO
Figure 4: Performance of different lifelong robot learning methods on the LIBERO benchmark.
5.2 Results and Analysis
Main Results.
Figure 3 presents the performance of the pretrained policy on both the original training tasks and zero-shot on unseen tasks, while Figure 4 summarizes three key metrics mentioned in Section 3.2 across all evaluated lifelong learning methods. Pretraining with FFT improves downstream transfer as expected, but still results in poor zero-shot performance, highlighting the need for more effective adaptation strategies. Besides, using frozen CLIP features is inefficient for robotic manipulation and performs even worse when pretraining is applied, likely due to overfitting in other sub-modules. Notably, there is a huge gap between multi-task (MT) and sequential fine-tuning (SeqFT). Among the full fine-tuning methods, ER achieves the best results, but this comes at the cost of storing and replaying large amounts of previous data (see Figure 1). LOTUS fails on the Object suite with the same code as on other suites, indicating poor robustness. Existing LoRA-based methods show lower forward transfer compared to FFT approaches, while DMPEL improves forward transfer by using dynamically synthesized experts, which reuse previous knowledge in a flexible way. Furthermore, our expert coefficient replay mechanism is highly effective in mitigating catastrophic forgetting, achieving near-zero NBT across all benchmarks without using oracle task identifiers. Conversely, L2M demonstrates weak forward transfer, likely due to its frozen action head. Although IsCiL reduces forgetting on simpler suites, it struggles with long-horizon tasks, presumably because of its fixed context encoding mechanism. In summary, DMPEL achieves efficient forward transfer and low catastrophic forgetting during lifelong learning, with only 1.2M trainable parameters (less than 0.7% of the policy) and minimal storage overhead.
((a))
((b))
((c))
Figure 5: Ablation studies. (a) Different algorithm settings on LIBERO-Goal (b) Different coefficient replay ratios on LIBERO-Object (c) Different expert synthesize intervals on LIBERO-Spatial.
Ablation Studies.
Figure 5(a) examines the effects of various algorithm settings. Utilizing the top-3 experts yields better performance compared to using only the top-1 expert or merging all experts. Additionally, orthogonal initialization of new experts helps to reduce interference. Besides, directly synthesizing policy parameters and assigning distinct coefficients to different submodules allows for more flexible and effective integration of knowledge. Figure 5(b) illustrates the impact of the coefficient replay (CR) ratio on overall performance. CR ratio refers to the proportion of router input-output pairs stored in the buffer for future replay. The results show that coefficient replay does not hinder forward transfer but helps mitigate forgetting efficiently, even when only a small fraction (e.g., 5%) of previous router input-output pairs is retained. Additionally, we demonstrate that replaying the context embeddings and coefficient vectors incurs a much lower storage overhead (approx. 6.7%) compared to replaying full expert demonstrations.
Figure 6: Visualization of expert coefficients on the LIBERO-Object benchmark.
Computational Overhead.
DMPEL introduces two additional computational stages: (1) encoding the current context with the frozen backbone and using the global router to compute routing coefficients, and (2) averaging the parameters of the top-$\delta$ experts to synthesize $\tilde{\boldsymbol{W}}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$, which incurs approximately $2\times\delta\times r\times(d_{\text{in}}+d_{\text{out}})$ FLOPs. Thereafter, the forward pass through the linear layer requires the usual $2\times d_{\text{in}}\times d_{\text{out}}$ FLOPs. In robotic manipulation, which typically spans hundreds of decision steps and whose observations change only gradually from one step to the next, we conjecture that re-synthesizing policy parameters every several steps can yield a favorable trade-off between computational cost and task performance. Figure 5(c) verifies that, as the expert synthesize interval increases from 1 to 5, the average inference time rapidly approaches that of the backbone-only baseline (approx. 30 ms), while the success rate suffers only a slight decline. The result also indicates that the expert router indeed dynamically selects low-rank experts tailored to the current context.
Visualization on Expert Coefficients.
We present a visualization of the expert coefficients for each sub-module during the lifelong learning process on the LIBERO-Object benchmark. As illustrated in Figure 6, significant knowledge sharing occurs across tasks, which contributes to the enhanced forward transfer capability of DMPEL. Notably, we observe that the low-rank expert learned for the action head in the first task is extensively reused in subsequent tasks, while the experts introduced later are rarely activated. This suggests that these inactive experts can be pruned to further improve storage efficiency without adversely affecting performance.
6 Conclusion
In this study, we introduce Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL) for lifelong robot learning. Our approach builds an low-rank expert library incrementally and utilizes a lightweight router to dynamically integrate these experts. This enables the robot to adapt flexibly to various scenarios over its lifetime. Additionally, by utilizing the modular nature of the policy, we address the issue of catastrophic forgetting through coefficient replay on the router, which allows for precise retrieval of experts on previous tasks. This method is significantly more efficient than the expensive demonstration replay on the entire policy. Our extensive experiments on the LIBERO benchmark show that DMPEL surpasses state-of-the-art lifelong learning baselines in terms of forward transfer and catastrophic forgetting while requiring only very few trainable parameters and storage.
Limitations.
We conduct all experiments using a rather small model and we only use in-domain data to do pretraining and adaptation. We plan to further scale up our policy or employ more advanced VLA models (over billions of parameters) as the backbone in the future. In addition, we seek to further analyze the mechanisms of knowledge transfer and to enhance lifelong learning algorithms with a solid theoretical foundation.
Appendix A Experimental Details
A.1 Policy and Algorithm Details
As illustrated in Section 4.1, we use an auto-regressive policy consisting of vision, text, and state encoders, an input modality fusion module, a temporal transformer, and an action head [40, 43]. Detailed hyperparameters of each submodule are summarized in Table 1.
We utilize the implementations of Experience Replay (ER) [22], Elastic Weight Consolidation (EWC) [27], and PackNet [30] as provided in LIBERO [43]. For ER, we employ a replay ratio of 20%. For EWC, we set the hyperparameters $\gamma=0.9$ and $\lambda=5\times 10^{4}$. For PackNet, we fix 25% of the parameters as important. When training with ER and EWC, we use two GPUs, each with a batch size of 16, ensuring the total batch size remains consistent with other methods. For additional details on these baselines, we refer readers to [43]. We directly use the official code of LOTUS [92]. Since LOTUS employs a distinct architecture, we adjust its parameters to be comparable to our policy and follow the same training pipeline. Additionally, LOTUS utilizes the DINO-v2 [14] foundation model instead of CLIP, as recommended by the ablation study in their paper.
For LoRA-based methods, we insert LoRA modules with rank $r=8$ into the 6-th to 11-th layers of the CLIP image and text encoders, and LoRA modules with $r=16$ into the temporal transformer. LoRA is applied to the query and value projection matrices. In TAIL and IsCiL, we learn task-specific fusion modules, state encoders, and action heads, as described in [40]. In contrast, these components are frozen in L2M following the original paper [41]. Due to the unavailability of publicly released code for TAIL, we implemented it based on the description in the paper [40]. We integrated the code for L2M and IsCiL into our codebase, as the original implementations operate with different policies across different benchmarks. For L2M, we use a pool size of 30 and the similarity loss coefficient $\lambda=0.5$. For IsCiL, we set the skill prototype $\mathcal{X}_{z}$ in $\mathcal{X}$ to be composed of 20 bases as in [42]. For DMPEL, we adopt the same LoRA settings in the image encoder, text encoder, and temporal transformer, and use LoRA experts with $r=16$ for all linear layers in other submodules. Besides, in the Long suite, we tune both the low-rank weights $\boldsymbol{A},\boldsymbol{B}$ and biases $\boldsymbol{b}$, whereas in the Goal, Spatial, and Object suites, we only tune the low-rank weights $\boldsymbol{A},\boldsymbol{B}$. However, both settings lead to fewer learnable parameters than TAIL [40], which learns these sub-modules separately for each task. We also employ dropout with probability $p=0.15$ on the coefficient for low-rank experts to avoid overfitting.
Table 1: Environment and Policy Architecture Configuration
Hyperparameter Value Hyperparameter Value
Vision & Text Encoder
Pretrained model CLIP ViT-B/16 [15] Image resolution 128*128
Proprioceptive States
Joint state dim. 7 Gripper state dim. 2
Modality Fusion
Number of hidden layers 1 Hidden size 256
Activation GELU
Temporal Transformer
Number of layers 6 Number of attn. heads 8
Embedding size 768 MLP hidden size 1024
Max sequence Length 10 Dropout 0.15
Action Head
Number of hidden layers 2 Hidden size 256
Number of Modes 5 Min. std 1e-4
Activation Softplus Action dim. 7
Expert router
Number of hidden layers 2 Hidden Activation GELU
Output Activation Sigmoid
A.2 Training Configuration
We generally follow the setup in the LIBERO benchmark for training and evaluation [43]. Detailed settings and hyperparameters are summarized in Table 2. For all experiments, we use either Nvidia A100 or H800 GPUs. We employ Distributed Data Parallel (DDP) for parallel training across 4 GPUs during pretraining, while using a single GPU for lifelong learning. We utilize the 16-bit floating point precision (FP16) training mode to accelerate the training process. To ensure reproducibility, we adopted three random seeds, 100, 200, 300, and we report the mean and standard deviation of the performance over three seeds throughout the paper.
Table 2: Training Hyperparameters
Hyperparameter Value Hyperparameter Value Hyperparameter Value
Optimizer
Training epochs 10 Batch size 32 Optimizer AdamW
Betas (0.9,0.999) Learning rate 1e-4 Anneal strategy cos
Weight decay 0.1 Gradient clip 100
Image Augmentation
Brightness 0.3 Saturation 0.3 Contrast 0.3
Hue 0.3 Color jitter prob. 0.9 Rotation 15
Translation 0.1 Affine prob. 0.6
Evaluation
Epochs per eval. 2 Eval. episodes 20
A.3 Benchmark Details
We present the language instructions for the tasks in four lifelong learning LIBERO task suites in Table 3. Although some tasks have similar descriptions, they are not identical due to differences in environment configurations, such as varying spatial layouts, objects, or goal positions. Each task provides 50 successful expert demonstrations for imitation learning.
Table 3: Taks Instructions of LIBERO Task Suites
Suite Instructions
open the middle drawer of the cabinet
put the bowl on the stove
put the wine bottle on top of the cabinet
open the top drawer and put the bowl inside
Goal put the bowl on top of the cabinet
push the plate to the front of the stove
put the cream cheese in the bowl
turn on the stove
put the bowl on the plate
put the wine bottle on the rack
pick up the black bowl between the plate and the ramekin and place it on the plate
pick up the black bowl next to the ramekin and place it on the plate
pick up the black bowl from table center and place it on the plate
pick up the black bowl on the cookie box and place it on the plate
Spatial pick up the black bowl in the top drawer of the wooden cabinet and place it on the plate
pick up the black bowl on the ramekin and place it on the plate
pick up the black bowl next to the cookie box and place it on the plate
pick up the black bowl on the stove and place it on the plate
pick up the black bowl next to the plate and place it on the plate
pick up the black bowl on the wooden cabinet and place it on the plate
pick up the alphabet soup and place it in the basket
pick up the cream cheese and place it in the basket
pick up the salad dressing and place it in the basket
pick up the bbq sauce and place it in the basket
Object pick up the ketchup and place it in the basket
pick up the tomato sauce and place it in the basket
pick up the butter and place it in the basket
pick up the milk and place it in the basket
pick up the chocolate pudding and place it in the basket
pick up the orange juice and place it in the basket
put both the alphabet soup and the tomato sauce in the basket
put both the cream cheese box and the butter in the basket
turn on the stove and put the moka pot on it
put the black bowl in the bottom drawer of the cabinet and close it
Long put the white mug on the left plate and put the yellow and white mug on the right plate
pick up the book and place it in the back compartment of the caddy
put the white mug on the plate and put the chocolate pudding to the right of the plate
put both the alphabet soup and the cream cheese box in the basket
put both moka pots on the stove
put the yellow and white mug in the microwave and close it
Appendix B Additional Results
B.1 Learning curves
We present the learning curves for all methods in four lifelong learning LIBERO task suites in Figure 7-10 . The $y$-axis indicates the success rate, while the $x$-axis depicts the agent’s training process over the course of lifelong learning. For example, the Task 1 subplot in the figure shows the agent’s performance on the first task as it learns ten tasks sequentially.
Figure 7: Visualization of learning curves on the LIBERO-Goal benchmark.
Figure 8: Visualization of learning curves on the LIBERO-Spatial benchmark.
Figure 9: Visualization of learning curves on the LIBERO-Object benchmark.
Figure 10: Visualization of learning curves on the LIBERO-Long benchmark.

### v2

@@
 To address these limitations, we propose Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL) for lifelong robot learning.
-DMPEL progressively learn a low-rank expert library and employs a lightweight router to dynamically combine experts into an end-to-end policy, facilitating flexible behavior during lifelong adaptation.
-Moreover, by leveraging the modular structure of the fine-tuned parameters, we introduce coefficient replay to guide the router in accurately retrieving frozen experts for previously encountered tasks, thereby mitigating catastrophic forgetting.
-This method is significantly more storage- and computationally-efficient than applying demonstration replay to the entire policy.
-Extensive experiments on the lifelong manipulation benchmark LIBERO demonstrate that our framework outperforms state-of-the-art lifelong learning methods in success rates across continual adaptation, while utilizing minimal trainable parameters and storage.
+DMPEL progressively builds a low-rank expert library and employs a lightweight router to dynamically combine experts into an end-to-end policy, enabling flexible and efficient lifelong forward transfer.
+Furthermore, by leveraging the modular structure of the fine-tuned parameters, we introduce expert coefficient replay, which guides the router to accurately retrieve frozen experts for previously encountered tasks.
+This technique mitigates forgetting while being significantly more storage- and computation-efficient than experience replay over the entire policy.
+Extensive experiments on the lifelong robot learning benchmark LIBERO demonstrate that our framework outperforms state-of-the-art lifelong learning methods in success rates during continual adaptation, while utilizing minimal trainable parameters and storage.
+Code is publicly released for reproducibility.
 1 Introduction
-A generalist agent should have the capability to learn and adapt continuously throughout its lifetime, usually termed as lifelong learning [1, 2, 3, 4, 5, 6].
+A generalist agent should have the capability to learn and adapt continuously throughout its lifetime, usually termed as lifelong learning (De Lange et al., 2021; Masana et al., 2022; Mendez & Eaton, 2023; Wang et al., 2024a).
 The longstanding challenges are enabling forward transfer (i.e., leveraging knowledge from previous tasks to quickly adapt to new ones) and avoiding catastrophic forgetting (i.e., retaining previously acquired knowledge when learning new tasks) under limited computational and memory capacity.
-Due to the notorious sample inefficiency of the traditional tabula rasa approach in robotics, researchers have recently explored the pretrain-then-finetune paradigm [7, 8, 9, 10, 11, 12] that originates from the vision and language domains [13, 14, 15, 16, 17, 18, 19, 20].
+Due to the notorious sample inefficiency of the traditional tabula rasa approach in robotics, researchers have recently explored the pretrain-then-finetune paradigm (Brohan et al., 2022; 2023; Kim et al., 2025; Black et al., 2024) that originates from the vision and language domains (Bommasani et al., 2021; Oquab et al., 2024; Radford et al., 2021; Brown et al., 2020).
 Specifically, this paradigm first pretrain a policy, often referred to as a vision-language-action (VLA) model, on large-scale datasets, and then fine-tune it for various downstream tasks.
+Figure 1: Existing lifelong learning methods: (a) Replay methods; (b) Regularization methods; (c) Architectural methods; (d) Performance on LIBERO-Object.
+TAIL with LoRA utilizes significantly fewer resources than ER/EWC with FFT and exhibits no forgetting when provided with a task identifier, but demonstrates lower forward transfer.
+In contrast, DMPEL leverages a low-rank expert library and coefficient replay (CR==5%) to achieve better forward transfer and near-zero forgetting.
 However, in the context of lifelong learning, naively applying full fine-tuning to sequentially arriving robotic tasks can result in severe forgetting and suboptimal performance, failing to meet our requirements.
 As depicted in Figure 1, prior work in lifelong robot learning can be broadly classified into three distinct approaches.
-Replay methods [8, 21, 22, 23, 24, 25] retain previous data (e.g., vision-language web data, robot pretraining data, and data from seen tasks) and mix it with in-domain data during training on new tasks.
+Replay methods (Brohan et al., 2023; Lopez-Paz & Ranzato, 2017; Chaudhry et al., 2019; Shin et al., 2017; Xie & Finn, 2022) retain previous data (e.g., vision-language web data, robot pretraining data, and data from seen tasks) and mix it with in-distribution data during training on new tasks.
 These methods leads to generalizable policies, but also necessitates huge storage space and computational overhead for retaining old samples.
-Regularization methods [26, 27, 28] balance between new and old tasks by restricting the update of parameters.
-For example, EWC [27] require an entire copy of frozen model from the previous task and entail substantial computation in estimating the importance of parameters.
+Regularization methods (Zenke et al., 2017; Kirkpatrick et al., 2017; Li & Hoiem, 2017) balance between new and old tasks by restricting the update of parameters.
+For example, EWC (Kirkpatrick et al., 2017) require an entire copy of frozen model from the previous task and entail substantial computation in estimating the importance of parameters.
 In addition, they usually struggle in long task sequence due to plasticity-stability dilemma.
-In contrast, architectural methods [29, 30, 31] address catastrophic forgetting by explicitly learning task-specific parameters.
-The widely adopted parameter-efficient fine-tuning (PEFT) techniques [32, 33], such as adapters [34, 35], low-rank adaptation (LoRA) [36], and prompt tuning [37, 38, 39], have proven highly effective for steering frozen foundation models by integrating small, learnable modules in single-task adaptation.
-In the lifelong learning setting, a practical solution, as exemplified by TAIL [40], involves initializing and training task-specific adapters for each task, subsequently freezing them, and retrieving the appropriate adapter based on the oracle task index.
+In contrast, architectural methods (Rusu et al., 2016; Mallya & Lazebnik, 2018; Mallya et al., 2018; Ge et al., 2024) address catastrophic forgetting by explicitly learning task-specific parameters.
+The widely adopted parameter-efficient fine-tuning (PEFT) techniques (Ding et al., 2023), such as adapters (Houlsby et al., 2019; Chen et al., 2022), low-rank adaptation (LoRA) (Hu et al., 2021), and prompt tuning (Jia et al., 2022; Li & Liang, 2021; Lester et al., 2021), have proven highly effective for steering frozen foundation models by integrating small, learnable modules in single-task adaptation.
+In the lifelong learning setting, a practical solution, as exemplified by TAIL (Liu et al., 2024b), involves initializing and training task-specific adapters for each task, subsequently freezing them, and retrieving the appropriate adapter based on the oracle task index.
 Such method allows efficient adaptation with only a small amount of task-specific parameters, but also exhibits certain limitations.
@@
 Second, failing to effectively share knowledge across isolated adapters result in poor forward transfer.
-Some prior arts construct a shared adapter pool and enable automatic timestep-level retrieval by, for example, performing query-key matching to select the one with highest similarity from a pool of adapters [41], or using multifaceted prototypes to represent the cluster center in the state space for each adapter [42], but usually suffer from inaccurate retrieval and suboptimal performance.
-Figure 1: Existing lifelong learning methods: (a) Replay methods; (b) Regularization methods; (c) Architectural methods; (d) Performance on LIBERO-Object.
-TAIL with LoRA utilizes significantly fewer resources than ER/EWC with FFT and exhibits no forgetting when provided with a task identifier, but demonstrates lower forward transfer.
-In contrast, DMPEL leverages a low-rank expert library and coefficient replay (CR$=$5%) to achieve better forward transfer and near-zero forgetting.
-In this paper, we introduce Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL) to overcome these challenges in lifelong robot learning.
-DMPEL incrementally builds a library of low-rank experts and uses a lightweight router to dynamically assemble these experts into a unified policy, enabling the agent to adapt flexibly throughout continual adaptation.
-By exploiting the modularity of fine-tuned parameters, DMPEL incorporates coefficient replay at the router, enabling the policy to accurately integrate experts for all previously seen tasks and thereby effectively preventing catastrophic forgetting.
-This approach is significantly more efficient in terms of storage and computation than replaying expert demonstrations for the entire policy.
-Through extensive evaluations on the LIBERO lifelong manipulation benchmark [43], we show that our method achieves higher forward transfer and lower catastrophic forgetting than existing baselines while requiring minimal trainable parameters and storage.
-Code is publicly released for reproducibility.
+Some prior arts construct a shared adapter pool and enable automatic timestep-level retrieval by, for example, performing query-key matching to select the one with highest similarity from a pool of adapters (Schmied et al., 2023), or using multifaceted prototypes to represent the cluster center in the state space for each adapter (Lee et al., 2024), but usually suffer from inaccurate retrieval and suboptimal performance.
+In this paper, we introduce the Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL) to address key challenges in lifelong robot learning.
+As illustrated in Figure 2, DMPEL features two core innovations: (1) it incrementally constructs a library of low-rank experts and employs a lightweight router to dynamically compose these experts into a unified policy based on the current context.
+This design allows the agent to flexibly adapt while leveraging all previously acquired parametric knowledge, thereby enhancing forward transfer; (2) it uses expert coefficient replay to regularize the router in accurately integrating experts for all previously encountered tasks, exploiting the modularity of fine-tuned parameters to prevent catastrophic forgetting.
+Notably, replaying low-dimensional embeddings and coefficients using the router is significantly more efficient than replaying demonstrations that require the entire policy in terms of storage and computation.
+Extensive evaluations on the lifelong manipulation benchmark LIBERO (Liu et al., 2024a) demonstrate that our method achieves superior forward transfer with reduced catastrophic forgetting compared to existing baselines, all while requiring minimal trainable parameters and storage.
+Figure 2: Overview of the proposed method DMPEL.
 2 Related Work
 Foundation Models for Robotics.
-The traditional tabula rasa paradigm in robotics is widely criticized for its sample inefficiency, prompting researchers to explore the use of foundation models to improve downstream task transfer [44, 45].
-One prominent direction involves transferring general visual representations: some studies directly employ foundation models from other domains, such as CLIP [15], SAM [18], and DINOv2 [14], as vision encoders [46, 47], while others pretrain visual representations using self-supervised learning [16, 17] on egocentric video datasets like R3M [48], MVP [49], and VC-1 [50].
-Another emerging line of research focuses on pretraining large-scale policies on diverse robotic tasks to capture generalizable behavior priors, for example, RT-1 [7], RT-2 [8], Gato [9], Octo [10], OpenVLA [11], and $\pi_{0}$ [12].
+The traditional tabula rasa paradigm in robotics is widely criticized for its sample inefficiency, prompting researchers to explore the use of foundation models to improve downstream task transfer (Firoozi et al., 2023; Hu et al., 2023).
+One prominent direction involves transferring general visual representations: some studies directly employ foundation models from other domains as vision encoders (Parisi et al., 2022; Yuan et al., 2022), such as CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2024), while others pretrain visual representations on egocentric video datasets (Nair et al., 2023; Majumdar et al., 2024).
+Another emerging line of research focuses on pretraining large-scale policies on diverse robotic tasks to capture generalizable behavior priors, for example, RT-2 (Brohan et al., 2023), OpenVLA (Kim et al., 2025), and $\pi_{0}$ (Black et al., 2024).
 However, when adapting these models to downstream tasks, most approaches either rely on the frozen pretrained model (i.e., zero-shot transfer) or require costly full fine-tuning.
-Some prior work has explored parameter-efficient fine-tuning (PEFT) by integrating task-specific modules into the pretrained vision encoder [51, 52, 53] or temporal transformer [54, 55, 56].
-Despite these advancements, the majority of these methods focus on single-stage adaptation, whereas our approach targets lifelong adaptation to a sequence of new tasks.
+Some prior work has explored parameter-efficient fine-tuning (PEFT) by integrating task-specific modules into the pretrained vision encoder (Sharma et al., 2023; Marza et al., 2024) or temporal transformer (Liang et al., 2022; Xu et al., 2023; Qiao et al., 2023).
+Despite these advancements, the majority of these arts focus on single-task adaptation, whereas our approach targets lifelong adaptation to a sequence of new tasks.
 Lifelong Learning with PEFT.
-Compared to full fine-tuning, PEFT optimizes only a small amount of inserted parameters and keeps the rest untouched, greatly reducing the storage cost of tuned parameters and mitigating the risk of overfitting in low-data regimes [32].
-Notably, PEFT techniques share conceptual similarities with architectural methods in lifelong learning.
-Most studies on lifelong learning with PEFT predominantly utilize prompt-based tuning [37, 39] to adapt frozen pretrained model for tasks such as image classification (e.g., L2P [57], DualPrompt [58], CODA-Prompt [59], and DAP [60]) and text classification (e.g., EPI [61], and Progressive Prompts [62]).
-Current research primarily investigates strategies for dynamically selecting or generating instance-specific prompts [63].
-Recognizing the representational limitations of prompt tuning, subsequent studies have extended to more powerful PEFT techniques, including LAE [64], HiDe-PET [65].
-Despite the demonstrated success of PEFT-based lifelong learning in classification tasks, its application remains largely unexplored in robotics.
-A few pioneering works employ LoRA [36] to store task-specific knowledge, yet diverge in their retrieval mechanisms: TAIL [40] relies on oracle task identifier, L2M [41] implements query-key matching between the current context (the state embedding from the frozen pretrained model) and learnable keys, while IsCiL [42] establishes multifaceted prototypes through K-means clustering in the state space, with input states retrieving the closest LoRA parameters based on proximity.
+Compared to full fine-tuning, PEFT optimizes only a small amount of inserted parameters while keeping the rest unchanged, which shares conceptual similarities with architectural methods in lifelong learning (Ding et al., 2023).
+Most studies on PEFT-based lifelong learning predominantly utilize prompt-based tuning (Jia et al., 2022; Lester et al., 2021) to adapt frozen pretrained model for tasks such as image classification, e.g., L2P (Wang et al., 2022c), DualPrompt (Wang et al., 2022b), CODA-Prompt (Smith et al., 2023), DAP (Jung et al., 2023), and text classification, e.g., EPI (Wang et al., 2023b), Progressive Prompts (Razdaibiedina et al., 2023).
+Current research primarily investigates strategies for dynamically selecting or generating instance-specific prompts (Zhou et al., 2024).
+Recognizing the representational limitations of prompt tuning, subsequent studies have extended to more powerful PEFT techniques, including LAE (Gao et al., 2023), HiDe-PET (Wang et al., 2025), and SKILL (Ge et al., 2024).
+Despite the success in incremental classification tasks, its application remains largely unexplored in robotics.
+A few pioneering works employ LoRA (Hu et al., 2021) to store task-specific knowledge, yet diverge in their retrieval mechanisms: TAIL (Liu et al., 2024b) relies on oracle task identifiers, L2M (Schmied et al., 2023) implements query-key matching between the current context and learnable keys, while IsCiL (Lee et al., 2024) establishes multifaceted prototypes through K-means clustering in the state space, with input states retrieving the closest LoRA parameters based on proximity.
 Our method incrementally constructs a parameter-efficient expert library and learns a lightweight router to dynamically retrieve and combine these experts.
 Mixture of Experts (MoE) and Model Fusion.
-The Mixture of Experts (MoE) approach, which involves training multiple specialized models for specific subtasks, has gained significant attention [66, 67, 68].
-A key method for combining these experts is parameter ensemble, supported by the linear mode connectivity phenomenon, which shows models converging to a single low-loss basin [69].
-Task Arithmetic [70] introduces task vectors, representing the difference between fine-tuned and pretrained models, enhancing multi-task model performance when merged.
-Follow-up work [71, 72] improves this by learning input-conditioned layer-wise fusion coefficients, increasing flexibility and performance.
-Model fusion techniques also apply to PEFT, where previously learned adapters are reused for efficient task adaptation [73, 74, 75, 76], including LoRAHub [77] and AdapterFusion [78].
-In image generation, merging separate LoRAs facilitates joint style and content generation [79, 80, 81].
-In robotics, models like MELA [82], MoE-Loco [83], MORE [84], and SDP [85] create versatile, adaptive behaviors for multiple tasks by fusing expert sets.
-Our method aims to enhance forward transfer in lifelong robot learning by flexibly fusing previously learned low-rank experts for each sub-module in the policy.
-Besides, we exploit the modularized design by replaying coefficient on the router to mitigate catastrophic forgetting.
+MoE approach involves training multiple specialized models for specific subtasks and has gained significant attention (Fedus et al., 2022; Muqeeth et al., 2024).
+A key method for combining these experts is parameter ensemble, supported by the linear mode connectivity phenomenon, which shows models converging to a single low-loss basin (Frankle et al., 2020).
+Task Arithmetic (Ilharco et al., 2023) introduces task vectors, representing the difference between fine-tuned and pretrained models, enhancing multi-task model performance when merged.
+Follow-up work (Yang et al., 2024; Tang et al., 2024) improves this by learning input-conditioned layer-wise fusion coefficients, increasing flexibility and performance.
+Model fusion techniques also apply to PEFT, where previously learned adapters are reused for efficient task adaptation (Wu et al., 2023; 2024; Wang et al., 2022a; Zhao et al., 2024), including LoRAHub (Huang et al., 2024) and AdapterFusion (Pfeiffer et al., 2021).
+In robotics, models like MELA (Yang et al., 2020), MoE-Loco (Huang et al., 2025), MORE (Zhao et al., 2025), and SDP (Wang et al., 2024b) create versatile, adaptive behaviors for multiple tasks by fusing expert sets.
+Our method aims to enhance forward transfer in lifelong robot learning by flexibly fusing previously learned low-rank experts for each sub-module in the policy according to the current context.
+Besides, we exploit the modularized design by replaying expert coefficient on the router to mitigate catastrophic forgetting.
 3 Problem Formulation
@@
 The ultimate objective of robot learning is to search for a policy $\pi$ that maximizes the expected success rate in reaching the goal: $\max_{\pi}J(\pi)=\mathbb{E}_{s_{t},a_{t}\sim\pi,d_{0}}\left[\sum_{t=1}^{\mathcal{H}}l(s_{t})\right]$.
-Due to the significant challenge of sparse reward in robot manipulation, we consider a more practical lifelong imitation learning setting [43, 40], where a robot sequentially learns over a stream of tasks $\{\mathcal{T}^{1},\cdots,\mathcal{T}^{K}\}$ given an expert demonstration dataset $\mathcal{D}^{k}=\{\tau_{n}^{k}\}_{n=1}^{N}$ for each task $\mathcal{T}^{k}=(d_{0}^{k},l^{k})$, $1\leq k\leq K$.
+Due to the significant challenge of sparse reward in robot manipulation, we consider a more practical lifelong imitation learning setting (Liu et al., 2024a; b), where a robot sequentially learns over a stream of tasks $\{\mathcal{T}^{1},\cdots,\mathcal{T}^{K}\}$ given an expert demonstration dataset $\mathcal{D}^{k}=\{\tau_{n}^{k}\}_{n=1}^{N}$ for each task $\mathcal{T}^{k}=(d_{0}^{k},l^{k})$, $1\leq k\leq K$.
 Each expert trajectory $\tau_{n}^{k}$ can be represented as $(o_{0},a_{0},\cdots,o_{\mathcal{H}})$, in which observation $o_{t}$ at each step $t$ includes RGB images and proprioceptive states.
 Following common pratice in partially observable MDPs, we approximate the current state $s_{t}$ by stacking prior observations, i.e., $s_{t}=o_{\leq t}=(o_{0},\cdots,o_{t})$.
-We perform behavior cloning [86] to learn a stochastic policy $\pi_{\theta}$ by minimizing the negative log-likelihood loss $L$:
+We perform behavior cloning (Bain & Sammut, 1995) to learn a stochastic policy $\pi_{\theta}$ by minimizing the negative log-likelihood loss $L$:
 $\min_{\theta}L(\theta)=-\mathbb{E}_{\tau_{n}^{k}\sim\mathcal{D}^{k}}\left[\sum_{t=0}^{\mathcal{H}-1}\log\pi_{\theta}\left(a_{t}|o_{\leq t},l^{k}\right)\right].$ (1)
 3.2 Evaluation Metrics
-We use three metrics to evaluate the performance of lifelong robot learning [43, 87]: forward transfer (FWT), negative backward transfer (NBT), area under the success rate curve (AUC).
+We use three metrics to evaluate the performance of lifelong robot learning (Liu et al., 2024a; Rodríguez et al., 2018): forward transfer (FWT), negative backward transfer (NBT), area under the success rate curve (AUC).
 Formally, we evaluate the policy on current task $\mathcal{T}^{k}$ when reaching a checkpoint $c\in\mathcal{C}$ (i.e., every multiple epochs of training) and we denote the success rate during evaluation as $s_{k,k,c}$.
@@
 This checkpoint $c_{k}^{*}$ will be further evaluated on previous tasks $\mathcal{T}^{j},j\leq k$, where the success rate is denoted as $s_{k,j}$.
-Then we can define the three metrics as:
+Intuitively, higher FWT indicates faster adaptation to new tasks, lower NBT indicates less catastrophic forgetting on old tasks, and higher AUC demonstrates an overall better performance across tasks and a good balance between FWT and NBT.
+These metrics can be formally defined as:
 FWT $=\frac{1}{K}\sum_{k=1}^{K}\left[\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}s_{k,k,c}\right],$ (2)
@@
 AUC $=\frac{1}{K}\sum_{k=1}^{K}\left[\frac{1}{K\!-\!k\!+\!1}\Big(\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}s_{k,k,c}+\!\!\sum_{l=k+1}^{K}\!\!s_{l,k}\Big)\right].$ (4)
-Intuitively, higher FWT indicates faster adaptation to new tasks, lower NBT indicates less catastrophic forgetting on old tasks, and higher AUC demonstrates an overall better performance across tasks and a good balance between FWT and NBT.
 4 Proposed Method
 4.1 Base Policy Architecture
-We use an autoregressive policy [11, 88] similar to the one used in the LIBERO benchmark [40, 43] and focus on developing better lifelong learning algorithms.
+We use a policy similar to the one used in the LIBERO benchmark (Liu et al., 2024a; b) and focus on developing better lifelong learning algorithms.
 The policy consists of vision, text, and state encoders, an input modality fusion module, a temporal transformer, and an action head.
@@
 In experiments, we use images from two cameras, the agent-view image and the eye-in-hand image.
-We employ the pretrained CLIP ViT-B/16 model [15] to serve as the vision encoder $\mathcal{E}_{I}$ and the text encoder $\mathcal{E}_{L}$.
+We employ the pretrained CLIP ViT-B/16 model (Radford et al., 2021) to serve as the vision encoder $\mathcal{E}_{I}$ and the text encoder $\mathcal{E}_{L}$.
 For the joint states and gripper states, we learn separate linear projection layers $\mathcal{E}_{P}$ to project the low-dimensional states to proprioceptive embeddings.
@@
 Modality fusion.
-We utilize a feature-wise linear modulation (FiLM) [89] to fuse language embeddings with the image and state embeddings.
+We utilize a feature-wise linear modulation (FiLM) (Perez et al., 2018) to fuse language embeddings with the image and state embeddings.
 Generally, given the original feature $x$, the conditional input $z$, and the fusion network $f_{\text{FiLM}}$, we can obtain the modulated feature $x^{\prime}=\gamma\odot x+\beta$ where $\gamma,\beta=f_{\text{FiLM}}(z)$.
@@
 Temporal transformer and action head.
-The causal temporal transformer backbone processes a sequence of multi-modal embeddings to produce a latent vector $g_{t}$ at each decision-making timestep.
-We employ a Gaussian Mixture Model (GMM)-based output head to compute the multi-modal distribution of manipulation actions [40, 43].
+The causal temporal transformer backbone $\mathcal{D}_{T}$ processes a sequence of multi-modal embeddings to produce a latent vector $g_{t}$ at each decision-making timestep.
+We employ a Gaussian Mixture Model (GMM)-based output head $\mathcal{D}_{H}$ to compute the multi-modal distribution of manipulation actions (Liu et al., 2024b; a).
 During training, we optimize the negative log-likelihood loss between the action distribution and the ground truth action.
 At inference time, the robot executes the policy by using the mean of the Gaussian distribution with the highest density for end effectors.
-Figure 2: Overview of the proposed method DMPEL.
 4.2 Low-rank Expert Library
 Following the pretrain-then-finetune paradigm, we pretrain the policy on a large-scale robotic dataset and freeze it during the lifelong adaptation to unseen downstream tasks.
-Low-Rank Adaptation (LoRA) [36].
-A representative parameter-efficient fine-tuning (PEFT) method that introduces learnable low-rank matrices $\boldsymbol{A}\in\mathbb{R}^{d_{\text{in}}\times r}$, $\boldsymbol{B}\in\mathbb{R}^{r\times d_{\text{out}}}$ and integrate in parallel with the frozen weight matrix $\boldsymbol{W}_{0}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ through addition, i.e., $\boldsymbol{W}_{0}+\boldsymbol{A}\boldsymbol{B}$.
+Low-Rank Adaptation (LoRA) (Hu et al., 2021) is a representative PEFT method that introduces learnable low-rank matrices $\bm{A}\in\mathbb{R}^{d_{\text{in}}\times r}$, $\bm{B}\in\mathbb{R}^{r\times d_{\text{out}}}$ and integrate in parallel with the frozen weight matrix $\bm{W}_{0}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ through addition, i.e., $\bm{W}_{0}+\bm{A}\bm{B}$.
 By leveraging the low-rank decomposition $r\ll\min\{d_{\text{in}},d_{\text{out}}\}$, LoRA substantially reduces the number of trainable parameters.
-Existing PEFT methods learn a task-specific low-rank adapter $\Delta\boldsymbol{W}_{k}=\boldsymbol{A}_{k}\boldsymbol{B}_{k}$ for each task $\mathcal{T}_{k}$ in lifelong robot learning [40, 41].
+Existing methods usually learn a task-specific low-rank adapter $\Delta\bm{W}_{k}=\bm{A}_{k}\bm{B}_{k}$ for each task $\mathcal{T}_{k}$ in lifelong robot learning (Liu et al., 2024b; Schmied et al., 2023).
 Such design requires task identifier during inference, lacks flexibility due to fixed capacity, and also hinders knowledge sharing across tasks.
-Prior arts in vision and language domain, e.g., LoRAHub [77] and SD-LORA [90], use a task-wise learnable weight vector to combine existing low-rank experts for better performance on the new task.
+Prior arts in vision and language domain, e.g., LoRAHub (Huang et al., 2024) and SD-LORA (Wu et al., 2025), use a task-wise learnable weight vector to combine existing low-rank experts for better performance on the new task.
 A key insight is that these low-rank experts captures complementary knowledge from a variety of specific tasks and can be leveraged to solve a new task.
 However, lifelong learning in robotic domain could be even more challenging.
-First, compared to image or text classification [1, 3, 5] that requires only single forward inference, robotic manipulation is a sequential decision-making process with long horizon.
+First, compared to image or text classification (De Lange et al., 2021; Masana et al., 2022; Wang et al., 2024a) that requires only single forward inference, robotic manipulation is a sequential decision-making process with long horizon.
 Besides, robot learning requires a mixed types of knowledge, for example, visual concepts, textual task goals, different spatial relationships, and successful adaptation to a new task requires a combination of adaptation in each sub-modules.
-Following this motivation, we propose to learn a progressive low-rank expert library on top of the pretrained policy and dynamically reuse low-rank experts from prior tasks $j\leq k-1$ to facilitate efficient forward transfer to new task $\mathcal{T}_{k}$.
-We equip each chosen pretrained linear layer =$\mathcal{F}=\{\boldsymbol{W}_{0},\boldsymbol{b}_{0}\}$ with a low-rank expert library =$\mathcal{L}=\{\boldsymbol{A}_{j},\boldsymbol{B}_{j},\boldsymbol{b}_{j}\}_{j=1}^{k}$ and use a global lightweight router $\mathcal{R}$ to dynamically integrate pretrained and task-specific knowledge based on the current context, providing a more flexible solution for adaptation.
-Specifically, the router $\mathcal{R}:\mathbb{R}^{d_{r}}\rightarrow\mathbb{R}^{6\times k}$ is implemented as a multi-layer perceptron (MLP) to obtain the dynamic coefficient for each low-rank expert.
-The router input $\boldsymbol{r}_{t}=[f^{v}_{t-T:t},f^{l},p_{t-T:t}]\in\mathbb{R}^{d_{r}}$, where $\cdot_{t-T:t}$ indicates mean pooling over the $T$-step context window.
+Following this motivation, we propose to learn a progressive low-rank expert library on top of the pretrained policy and dynamically reuse low-rank experts from prior tasks to facilitate efficient forward transfer to new task $\mathcal{T}_{k}$.
+We equip each chosen pretrained linear layer $\mathcal{F}=\{\bm{W}_{0},\bm{b}_{0}\}$ with a low-rank expert library $\mathcal{L}=\{\bm{A}_{j},\bm{B}_{j},\bm{b}_{j}\}_{j=1}^{k}$ and use a lightweight router $\mathcal{R}$ to dynamically integrate pretrained and task-specific knowledge based on the current context, providing a more flexible solution for adaptation.
+Specifically, the router $\mathcal{R}:\mathbb{R}^{d_{r}}\rightarrow\mathbb{R}^{M\times k}$ is implemented as a multi-layer perceptron (MLP) to obtain the dynamic coefficient for each low-rank expert.
+The router input $\bm{r}_{t-T:t}=[\tilde{f}^{v}_{t-T:t},\tilde{f}^{l}_{t-T:t},p_{t-T:t}]\in\mathbb{R}^{d_{r}}$, where $\cdot_{t-T:t}$ indicates mean pooling over the $T$-step context window.
 This aggregated representation contains information streams from the visual embeddings, the task instruction embeddings, and also the low-dimensional proprioceptive states, and hence comprehensively encodes the current context.
-Then we can dynamically obtain the expert coefficient:
-$\boldsymbol{c}_{t}=[\boldsymbol{c}^{v}_{t};\boldsymbol{c}^{l}_{t};\boldsymbol{c}^{s}_{t};\boldsymbol{c}^{f}_{t};\boldsymbol{c}^{t}_{t};\boldsymbol{c}^{h}_{t}]=\text{top-k}(\mathcal{R}(\boldsymbol{r}_{t}),\delta),\quad\boldsymbol{c}_{t}^{\cdot}=[c_{1,t}^{\cdot},c_{2,t}^{\cdot},\cdots,c_{k,t}^{\cdot}]\in\mathbb{R}^{k}.$ (5)
-[…]
-$\boldsymbol{y}=\tilde{\boldsymbol{W}}\boldsymbol{x}+\tilde{\boldsymbol{b}}=\underbrace{\boldsymbol{W}_{0}\boldsymbol{x}+\boldsymbol{b}_{0}}_{\text{pretrained}}+\underbrace{\sum_{j=1}^{k}c_{j,t}\boldsymbol{A}_{j}\sum_{j=1}^{k}c_{j,t}\boldsymbol{B}_{j}\boldsymbol{x}+\sum_{j=1}^{k}c_{j,t}\boldsymbol{b}_{j}}_{\text{fine-tuned}}.$ (6)
+Then we can obtain the expert coefficient:
+$\bm{c}_{t}=\text{top-k}(\mathcal{R}(\bm{r}_{t-T:t}),\delta)\in\mathbb{R}^{M\times k}.$ (5)
+Note that the output $\bm{c_{t}}$ is divided into $M$ distinct expert coefficient vectors $\bm{c}_{t}^{\diamond}=[c_{1,t}^{\diamond},c_{2,t}^{\diamond},\cdots,c_{k,t}^{\diamond}]\in\mathbb{R}^{k}$, each assigned to a specific sub-module $\diamond$ in the policy.
+In our case, we have $M=6$, consisting of vision encoder, text encoder, state encoder, modality fusion module, temporal transformer, and action head.
+This design allows each sub-module to use a different coefficient vector to integrate knowledge, providing larger flexibility in adaptation.
+The $\text{top-k}(\cdot,\delta)$ operator retains only the top $\delta$ over $k$ entries of the input vector at their original values, while setting all other coefficients to zero.
+This ensures efficient utilization of experts by reducing computational overhead while maintaining flexibility for adaptation.
+Finally, the input-conditioned linear layer $\{\tilde{\bm{W}},\tilde{\bm{b}}\}$ is the dynamic mixture of the pretrained weight $\mathcal{F}$ and the low-rank experts from the library $\mathcal{L}$:
+$\bm{y}$ $=\bm{x}\tilde{\bm{W}}+\tilde{\bm{b}}=\underbrace{\bm{x}\bm{W}_{0}+\bm{b}_{0}}_{\text{pretrained}}+\underbrace{\bm{x}\sum_{j=1}^{k}c_{j,t}\bm{A}_{j}\sum_{j=1}^{k}c_{j,t}\bm{B}_{j}+\sum_{j=1}^{k}c_{j,t}\bm{b}_{j}}_{\text{fine-tuned}}.$ (6)
 Upon the completion of task $\mathcal{T}_{k}$, we freeze all low-rank experts in the library as acquired knowledge and initialize new learnable expert for task $\mathcal{T}_{k+1}$.
-In order to reduce the interference between tasks, we employ the Gram-Schmidt orthogonalization method on $\boldsymbol{A}_{k+1}$ [59, 91], while other elements $\boldsymbol{B}_{k+1}$ and $\boldsymbol{b}_{k+1}$ are simply zero-initialized.
+In order to reduce the interference between tasks, we employ the Gram-Schmidt orthogonalization method on $\bm{A}_{k+1}$ (Smith et al., 2023; Wang et al., 2023a), while other elements $\bm{B}_{k+1}$ and $\bm{b}_{k+1}$ are simply zero-initialized.
 4.3 Expert Coefficient Replay
-To mitigate catastrophic forgetting, we freeze the learned low-rank expert module after each task’s completion for future retrieval.
+To mitigate catastrophic forgetting, we freeze the learned low-rank expert after each task’s completion for future retrieval.
 However, since the router is continuously updated on sequentially arriving tasks, incorrect integration of the expert, i.e., the expert coefficients are inaccurate, may lead to severe policy distortion and suboptimal performance.
-Inspired by replay methods [22, 24, 76], we propose an expert coefficient replay mechanism specifically designed for the lightweight router.
-The implementation involves two phases: During task finalization, we archive a subset of the router’s input-output pairs, i.e., the context embedding $\boldsymbol{r}$ and the coefficient vector $\boldsymbol{c}$, in a dedicated buffer $\mathcal{B}_{\text{CR}}$.
-When adapting to new tasks, the core idea is to enforce consistency by requiring the router to generate coefficients that closely match the stored historical coefficients given the same context.
-This is achieved through minimizing a mean squared error (MSE) objective:
-$L_{\text{CR}}(\mathcal{R})=\mathbb{E}_{{(\boldsymbol{r}_{i},\boldsymbol{c}_{i,\text{old}})\sim\mathcal{B}_{\text{CR}}}}\frac{1}{2}\left(\mathcal{R}(\boldsymbol{r}_{i})-\boldsymbol{c}_{i,\text{old}}\right)^{2}.$ (7)
-Compared with conventional expert demonstration replay that requires computation through the entire policy, our coefficient replay strategy demonstrates superior efficiency by focusing exclusively on the lightweight router and its inputs and outputs, yielding substantial reductions in both computational overhead and storage requirements.
+Inspired by replay methods (Chaudhry et al., 2019; Xie & Finn, 2022; Zhao et al., 2024), we propose an expert coefficient replay mechanism to regularize the lightweight router on previously encountered tasks.
+The implementation involves two phases: During task finalization, we archive a subset of the router’s input-output pairs, i.e., the context embedding $\bm{r}$ and the coefficient vector $\bm{c}$, in a dedicated buffer $\mathcal{B}_{\text{CR}}$.
+When adapting to new tasks, the core idea is to enforce consistency by requiring the router to generate coefficients that closely match the stored historical coefficients given the same context, which can be achieved through minimizing a mean squared error (MSE) objective:
+$L_{\text{CR}}(\mathcal{R})=\mathbb{E}_{{(\bm{r}_{i},\bm{c}_{i,\text{old}})\sim\mathcal{B}_{\text{CR}}}}\frac{1}{2}\left(\mathcal{R}(\bm{r}_{i})-\bm{c}_{i,\text{old}}\right)^{2}.$ (7)
+Compared to conventional experience replay methods that require storing entire trajectories and processing them through the entire policy, our expert coefficient replay strategy leverages the modularity of the low-rank expert library and focuses on regularizing the lightweight router using low-dimensional embeddings and coefficients, resulting in significant reductions in both computational overhead and storage requirements.
 5 Experiments
 5.1 Experimental Setup
+We present an overview of the experimental setup below, with details included in Appendix A.
 Benchmark.
-We conduct evaluations in LIBERO [43], a lifelong learning benchmark on robot manipulation tasks.
-The robot in the benchmark is situated in a tabletop environment and equipped with a 6-DOF arm and a parallel gripper.
+We conduct extensive evaluation in a lifelong robot learning benchmark LIBERO (Liu et al., 2024a).
+The robot is situated in a tabletop environment and equipped with a 6-DOF arm and a parallel gripper.
 We use four lifelong learning task suites: Goal, Spatial, Object, and Long.
-Each suite consists of 10 robotic manipulation tasks designed to investigate the transfer of knowledge related to task goals, spatial information, and various objects.
-Details can be found in Appendix A.3.
+Each suite consists of 10 sequentially arriving robotic manipulation tasks designed to investigate the transfer of knowledge related to task goals, spatial information, and various objects.
 The benchmark also includes LIBERO-90, a collection of 90 short-horizon diverse tasks that serves as a pretraining dataset for downstream transfer.
+All tasks are consistent with the formulation in Section 3, i.e., generating continuous actions to control the robot based on multi-modal input.
 Baselines.
-We consider a naive sequential fine-tuning baseline, three representative lifelong learning methods that involve tuning the entire policy, and several state-of-the-art LoRA-based methods:
-  • Sequential Fine-tuning (SeqFT) [43]: Naively perform full fine-tuning (FFT) or use frozen pretrained feature (FPF), i.e., freeze the spatial encoders and fine-tune the rest, on sequentially arriving robot learning tasks.
-  • Experience Replay (ER) [22], a replay method that maintains a memory buffer with samples from previous tasks.
-We store a subset of data after each task and uniformly sample a fixed number of replay data from the buffer along with new task data during training.
-  • Elastic Weight Consolidation (EWC) [27], a regularization method that uses the Fisher Information Matrix (FIM) to quantify the importance of each parameter and adds a regularization term to the learning objective to constrain network updates.
-  • PackNet [30], an architectural method that first trains and prunes the network to retain a fixed proportion of the most important parameters.
-The selected parameters are then fine-tuned and frozen to prevent forgetting.
-  • LifelOng knowledge Transfer Using Skills (LOTUS) [92] involves two stages: extracting temporal segment features by leveraging frozen foundation models for skill discovery, and hierarchical imitation learning with the meta-controller and a growing skill library.
-  • Task-specific Adapters for Imitation Learning (TAIL) [40] introduces separate LoRAs into the spatial encoder and temporal decoder for each new task.
-The fusion module, state encoder, and action head are also full fine-tuned and stored as task-specific adapters.
-TAIL requires an oracle task identifier to retrieve the corresponding adapters during inference.
-In order to align with other methods, we also provide the result of SeqFT (LoRA), which fine-tunes LoRA parameters sequentially.
-  • Learning to Modulate (L2M) [41] maintains a learnable modulation pool with keys and associated modulators.
-The embedded history of state tokens serves as the query vector for query-key matching, which finally steers the pretrained policy’s behavior.
-We follow the original paper to use a frozen fusion module and action head, and insert LoRA pools in the CLIP encoders and the temporal transformer backbone.
-  • Incremental Learning of Retrievable Skills for Continual Imitation Learning (IsCiL) [42] establishes multifaceted prototypes through K-means clustering in the state space, where proper LoRA-based skills is retrieved upon current input state.
-We define the task-specific skill the same as in TAIL [40], i.e., LoRA matrices for transformer, the fusion module, the state encoder, and the action head.
+We consider sequential fine-tuning (SeqFT) baselines (Liu et al., 2024a) that either perform full fine-tuning (FFT) or use frozen pretrained feature (FPF) naively on sequentially arriving tasks.
+We also consider three representative lifelong learning methods that fine-tune all parameters, including Experience Replay (ER) (Chaudhry et al., 2019), Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), and PackNet (Mallya & Lazebnik, 2018).
+We also include a hierarchical baseline, LifelOng knowledge Transfer Using Skills (LOTUS) (Wan et al., 2024).
+Besides, we consider several LoRA-based methods that use different low-rank expert inserting and retrieving mechanism, including Task-specific Adapters for Imitation Learning (TAIL) (Liu et al., 2024b), Learning to Modulate (L2M) (Schmied et al., 2023), and Incremental Learning of Retrievable Skills for Continual Imitation Learning (IsCiL) (Lee et al., 2024).
+Since TAIL requires oracle task identifiers, we also provide the result of SeqFT (LoRA) to align with other methods.
 Implementation Details.
-We utilize the open-source CLIP ViT/B-16 model [15] as vision and text encoders, while all other components of the policy is learned from robot demonstrations.
-We use a six-layer temporal transformer, using the same embedding size as the CLIP model $d=768$.
-The LoRA rank is set to 8 for CLIP encoders and 16 for others, the same as in [40].
+We utilize CLIP (Radford et al., 2021) as vision and text encoders, while all other components of the policy is learned from robot demonstrations.
+The LoRA rank is set to 8 for CLIP encoders and 16 for others, the same as in (Liu et al., 2024b).
 We pretrain the base policy on the LIBERO-90 dataset and use it as the starting point of lifelong learning.
@@
 All results are averaged over three random seeds.
-More details can be found in Appendix A.
-Figure 3: Performance of the pretrained policy on LIBERO
 Figure 3: Performance of different lifelong robot learning methods on the LIBERO benchmark.
 5.2 Results and Analysis
+Figure 4: Performance of the pretrained policy on LIBERO
 Main Results.
@@
 Utilizing the top-3 experts yields better performance compared to using only the top-1 expert or merging all experts.
-Additionally, orthogonal initialization of new experts helps to reduce interference.
-Besides, directly synthesizing policy parameters and assigning distinct coefficients to different submodules allows for more flexible and effective integration of knowledge.
+Besides, assigning distinct coefficients to different submodules allows for more flexible and effective integration of knowledge.
 Figure 5(b) illustrates the impact of the coefficient replay (CR) ratio on overall performance.
@@
 6.7%) compared to replaying full expert demonstrations.
-Figure 6: Visualization of expert coefficients on the LIBERO-Object benchmark.
+((a))
+((b))
+Figure 6: Analysis on expert coefficients. (a) Visualization of coefficients for each sub-module in the policy on LIBERO-Object (b) Influence on success rate (SR) when pruning underactivated experts.
 Computational Overhead.
-DMPEL introduces two additional computational stages: (1) encoding the current context with the frozen backbone and using the global router to compute routing coefficients, and (2) averaging the parameters of the top-$\delta$ experts to synthesize $\tilde{\boldsymbol{W}}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$, which incurs approximately $2\times\delta\times r\times(d_{\text{in}}+d_{\text{out}})$ FLOPs.
+DMPEL introduces two additional stages: (1) encoding the current context with the frozen backbone and using the global router to compute routing coefficients, and (2) averaging the parameters of the top-$\delta$ experts to synthesize $\tilde{\bm{W}}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$, which incurs approximately $2\times\delta\times r\times(d_{\text{in}}+d_{\text{out}})$ FLOPs.
 Thereafter, the forward pass through the linear layer requires the usual $2\times d_{\text{in}}\times d_{\text{out}}$ FLOPs.
@@
 30 ms), while the success rate suffers only a slight decline.
-The result also indicates that the expert router indeed dynamically selects low-rank experts tailored to the current context.
-Visualization on Expert Coefficients.
+The result also indicates that the router indeed dynamically selects low-rank experts tailored to the current context.
+Figure 7: t-SNE visualization of embeddings from the final block when evaluated on Task 2 from LIBERO-Object.
+Visualization Analysis.
 We present a visualization of the expert coefficients for each sub-module during the lifelong learning process on the LIBERO-Object benchmark.
-As illustrated in Figure 6, significant knowledge sharing occurs across tasks, which contributes to the enhanced forward transfer capability of DMPEL.
+As illustrated in Figure 6(a), significant knowledge sharing occurs across tasks, which contributes to the enhanced forward transfer capability of DMPEL.
 Notably, we observe that the low-rank expert learned for the action head in the first task is extensively reused in subsequent tasks, while the experts introduced later are rarely activated.
-This suggests that these inactive experts can be pruned to further improve storage efficiency without adversely affecting performance.
-6 Conclusion
-In this study, we introduce Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL) for lifelong robot learning.
-Our approach builds an low-rank expert library incrementally and utilizes a lightweight router to dynamically integrate these experts.
-This enables the robot to adapt flexibly to various scenarios over its lifetime.
-Additionally, by utilizing the modular nature of the policy, we address the issue of catastrophic forgetting through coefficient replay on the router, which allows for precise retrieval of experts on previous tasks.
-This method is significantly more efficient than the expensive demonstration replay on the entire policy.
-Our extensive experiments on the LIBERO benchmark show that DMPEL surpasses state-of-the-art lifelong learning baselines in terms of forward transfer and catastrophic forgetting while requiring only very few trainable parameters and storage.
-Limitations.
-We conduct all experiments using a rather small model and we only use in-domain data to do pretraining and adaptation.
-We plan to further scale up our policy or employ more advanced VLA models (over billions of parameters) as the backbone in the future.
-In addition, we seek to further analyze the mechanisms of knowledge transfer and to enhance lifelong learning algorithms with a solid theoretical foundation.
-Appendix A Experimental Details
+This suggests that these inactive experts can be pruned to further improve storage efficiency without adversely affecting performance, as shown in Figure 6(b).
+We also present t-SNE visualization in Figure 7, illustrating the latent embeddings from the temporal transformer’s final block during the evaluation on Task 2.
+DMPEL maintains embedding consistency after adapting to new tasks, whereas SeqFT results in substantial representation drift and catastrophic forgetting.
+Due to page limit, we provide additional results in Appendix B, including further ablation studies, visualization of trajectories, and lifelong learning curves.
+6 Conclusion and Limitation
+We present Dynamic Mixture of Progressive Parameter-Efficient Expert Library (DMPEL), a lifelong robot learning framework that incrementally builds a low-rank expert library and uses a lightweight router with coefficient replay to dynamically combine experts, achieving high forward transfer and low catastrophic forgetting with high efficiency in both storage and computation.
+Experiments on the LIBERO benchmark demonstrate that DMPEL achieves superior performance compared to state-of-the-art baselines.
+However, currently all experiments employ a rather small model with in-domain data for pretraining and adaptation.
+Future work will explore scaling up to larger and advanced policies (e.g., VLA models with over billions of parameters) and transferring to real robots, as well as deepen the theoretical understanding of knowledge transfer and retention in lifelong learning.
+Reproducibility Statement
+Our work is reproducible using the code available at https://github.com/HarryLui98/DMPEL. Additionally, we include detailed pseudocode and hyperparameter settings in the Appendix.
+Appendix A Experimental Setup
 A.1 Policy and Algorithm Details
-As illustrated in Section 4.1, we use an auto-regressive policy consisting of vision, text, and state encoders, an input modality fusion module, a temporal transformer, and an action head [40, 43].
-Detailed hyperparameters of each submodule are summarized in Table 1.
-We utilize the implementations of Experience Replay (ER) [22], Elastic Weight Consolidation (EWC) [27], and PackNet [30] as provided in LIBERO [43].
-For ER, we employ a replay ratio of 20%.
-For EWC, we set the hyperparameters $\gamma=0.9$ and $\lambda=5\times 10^{4}$.
-For PackNet, we fix 25% of the parameters as important.
+Policy Architecture.
+We use a policy consisting of vision, text, and state encoders, an input modality fusion module, a temporal transformer, and an action head (Liu et al., 2024b; a).
+We utilize the open-source CLIP ViT/B-16 model (Radford et al., 2021) as vision and text encoders, while all other components of the policy is learned from robot demonstrations.
+Detailed hyperparameters of each sub-module are summarized in Table 2, most of them are naturally inherited from the open-source LIBERO benchmark and CLIP model.
+Baselines.
+We consider naive sequential fine-tuning baselines, representative lifelong learning methods that involve tuning the entire policy, and several LoRA-based methods:
+  • Sequential Fine-tuning (SeqFT) (Liu et al., 2024a): Naively perform full fine-tuning (FFT) or use frozen pretrained feature (FPF), i.e., freeze the spatial encoders and fine-tune the rest, on sequentially arriving tasks.
+  • Experience Replay (ER) (Chaudhry et al., 2019), a replay method that maintains a memory buffer with samples from previous tasks.
+We store a subset of data after each task (20% in our experiments) and uniformly sample a fixed number of replay data from the buffer along with new task data during training.
+We use a batch size of 32 for old data, consistent with the batch size for new data, and mix them up for every training iteration.
+  • Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), a regularization method that uses the Fisher Information Matrix (FIM) to quantify the importance of each parameter and adds a regularization term to the learning objective to constrain network updates.
+We set the hyperparameters $\gamma=0.9$ and $\lambda=5\times 10^{4}$.
+  • PackNet (Mallya & Lazebnik, 2018), an architectural method that first trains and prunes the network to retain a fixed proportion (25% in our experiments) of the most important parameters.
+The selected parameters are then fine-tuned and frozen to prevent forgetting.
+  • LifelOng knowledge Transfer Using Skills (LOTUS) (Wan et al., 2024) involves two stages: extracting temporal segment features by leveraging frozen foundation models for skill discovery, and hierarchical imitation learning with the meta-controller and a growing skill library.
+Since LOTUS employs a distinct architecture, we adjust its parameters to be comparable to our policy and follow the same training pipeline.
+Additionally, LOTUS utilizes the DINO-v2 (Oquab et al., 2024) foundation model instead of CLIP, as recommended by the ablation study in their paper.
+  • Task-specific Adapters for Imitation Learning (TAIL) (Liu et al., 2024b) introduces separate LoRAs into the spatial encoder and temporal decoder for each new task.
+The fusion module, state encoder, and action head are also full fine-tuned and stored as task-specific adapters.
+TAIL requires an oracle task identifier to retrieve the corresponding adapters during inference.
+In order to align with other methods, we additionally provide the result of SeqFT (LoRA), which fine-tunes LoRA parameters sequentially.
+  • Learning to Modulate (L2M) (Schmied et al., 2023) maintains a learnable modulation pool with keys and associated modulators.
+The embedded history of state tokens serves as the query vector for query-key matching, which finally steers the pretrained policy’s behavior.
+We follow the original paper to use a frozen fusion module and action head, and insert LoRA pools in the CLIP encoders and the temporal transformer backbone.
+We use a pool size of 30 and the similarity loss coefficient $\lambda=0.5$.
+  • Incremental Learning of Retrievable Skills for Continual Imitation Learning (IsCiL) (Lee et al., 2024) establishes multifaceted prototypes through K-means clustering in the state space, where proper LoRA-based skills is retrieved upon current input state.
+We define the task-specific skill the same as in TAIL (Liu et al., 2024b), i.e., LoRA matrices for transformer, the fusion module, the state encoder, and the action head.
+We use CLIP encoded features as context embeddings and set the skill prototype $\mathcal{X}_{z}$ in $\mathcal{X}$ to be composed of 20 bases.
+Table 1: Policy Configurations
+Hyperparameter Value
+Vision & Text Encoder
+Pretrained model CLIP ViT-B/16
+Image resolution 128*128
+Proprioceptive States
+Joint state dim. 7
+Gripper state dim. 2
+Modality Fusion
+Number of hidden layers 1
+Hidden size 256
+Activation GELU
+Temporal Transformer
+Number of layers 6
+Number of attn. heads 8
+Embedding size 768
+MLP hidden size 1024
+Max sequence Length 10
+Dropout 0.15
+Action Head
+Number of hidden layers 2
+Hidden size 256
+Number of Modes 5
+Min. std 1e-4
+Activation Softplus
+Action dim. 7
+Expert router
+Number of hidden layers 2
+Hidden Activation GELU
+Output Activation Sigmoid
+Table 2: Training Hyperparameters
+Hyperparameter Value
+Optimizer
+Training epochs 10
+Batch size 32
+Optimizer AdamW
+Betas (0.9,0.999)
+Learning rate 1e-4
+Anneal strategy cos
+Weight decay 0.1
+Gradient clip 100
+Image Augmentation
+Brightness 0.3
+Saturation 0.3
+Contrast 0.3
+Hue 0.3
+Color jitter prob. 0.9
+Rotation 15
+Translation 0.1
+Affine prob. 0.6
+Evaluation
+Epochs per eval. 2
+Eval. episodes 20
+Algorithm 1 Training Process of DMPEL
+0: Pretrained policy $\pi_{\theta}$, router $\mathcal{R}$, low-rank expert library $\mathcal{L}=\varnothing$ for each layer $\mathcal{F}$, coefficient replay buffer $\mathcal{B}_{\text{CR}}=\varnothing$
+1:for $k=1,2,\cdots,K$ do
+2:  for each $\{\mathcal{F},\mathcal{L}\}$ do
+3:   Initialize $\bm{A}_{k}$, $\bm{B}_{k}$, (and optionally $\bm{b}_{k}$)
+4:   $\mathcal{L}\leftarrow\text{sg}(\mathcal{L})\cup\{\bm{A}_{k},\bm{B}_{k},\bm{b}_{k}\}$ # $\text{sg}(\cdot)$ means stop gradient, i.e., freezing previously learned low-rank experts
+5:  end for
+6:  for epochs $c=1,2\cdots,C$ do
+7:   for each mini-batch data from dataset $\mathcal{D}^{k}$ do
+8:    Compute the context embedding $\bm{r}$ using the frozen encoders in the pretrained policy $\pi_{\theta}$
+9:    Compute the expert coefficient vector $\bm{c}$ with the router $\mathcal{R}$ using Eq. (5)
+10:    for each $\{\mathcal{F},\mathcal{L}\}$ do
+11:     Obtain the expert coefficient $\bm{c}^{\diamond}$ according to which sub-module $\diamond$ the layer belongs to
+12:     Synthesize parameters $\tilde{\bm{W}},\tilde{\bm{b}}$ using Eq. (6)
+13:    end for
+14:    Compute behavior cloning loss $L$ using Eq. (1)
+15:    Update router parameters $\mathcal{R}$ and low-rank expert parameters $\bm{A}_{k}$, $\bm{B}_{k}$, (and optionally $\bm{b}_{k}$) to minimize $L$
+16:    if $c=C$ then $\mathcal{B}_{\text{CR}}\leftarrow\mathcal{B}_{\text{CR}}\cup\{(\bm{r},\bm{c})\}$ with a probability equals to the coefficient replay ratio
+17:   end for
+18:   if $c\in\mathcal{C}$ then Evaluate on task $\mathcal{T}^{k}$ to obtain success rate $s_{k,k,c}$ (Algorithm 2)
+19:  end for
+20:  for epochs $c=1,2\cdots,C$ do
+21:   for each mini-batch data from the buffer $\mathcal{B}_{\text{CR}}$ do
+22:    Compute expert coefficient replay loss $L_{\text{CR}}$ using Eq. (7)
+23:    Update router parameters $\mathcal{R}$ to minimize $L_{\text{CR}}$
+24:   end for
+25:  end for
+26:  for $j=1,2,\cdots,k$ do
+27:   Evaluate on task $\mathcal{T}^{j}$ to obtain success rate $s_{k,j}$ (Algorithm 2)
+28:  end for
+29:end for
+30: Compute metrics FWT, NBT, AUC using Eq. (2)-(4)
+Algorithm 2 Inference Process of DMPEL
+0: Pretrained policy $\pi_{\theta}$, router $\mathcal{R}$, each layer with pretrained weight and low-rank expert library $\{\mathcal{F},\mathcal{L}\}$, FIFO queues $F_{q}=\varnothing$, $R_{q}=\varnothing$ with maximum length $T$
+1:for $t=1,2,\cdots,\mathcal{H}$ do
+2:  Observe $o_{t}=(I^{1}_{t},\cdots,I^{N_{c}}_{t},p_{t})$ and language instruction $l$
+3:  if $T_{\text{syn}}\mid t$ then
+4:   Compute visual and textual embeddings $\tilde{f}^{v}_{t}$, $\tilde{f}^{l}_{t}$ using the frozen CLIP encoders $\mathcal{E}_{I}$, $\mathcal{E}_{L}$ ($\mathcal{F}$ only)
+5:   Compute the context embedding $\bm{r}_{t}=[\tilde{f}^{v}_{t},\tilde{f}^{l}_{t},p_{t}]$
+6:   $R_{q}\leftarrow R_{q}\cup\{\bm{r}_{t}\}$
+7:   Compute the expert coefficient vector $\bm{c}_{t}=\text{top-k}(\mathcal{R}(\bm{r}_{t-T:t}),\delta)$
+8:   for each $\{\mathcal{F},\mathcal{L}\}$ do
+9:    Obtain the expert coefficient $\bm{c}^{\diamond}_{t}$ according to which sub-module $\diamond$ the layer belongs to
+10:    Synthesize parameters $\tilde{\bm{W}},\tilde{\bm{b}}$ using Eq. (6)
+11:   end for
+12:  else
+13:   Reuse previous visual and text embeddings $\tilde{f}^{v}_{t}\leftarrow\tilde{f}^{v}_{t-1}$, $\tilde{f}^{l}_{t}\leftarrow\tilde{f}^{l}_{t-1}$
+14:   $R_{q}\leftarrow R_{q}\cup\{[\tilde{f}^{v}_{t}$, $\tilde{f}^{l}_{t},p_{t}]\}$
+15:  end if
+16:  Compute visual, textual, and state embeddings $f^{v}_{t}$, $f^{l}_{t}$, $f^{s}_{t}$ using encoders $\mathcal{E}_{I}$, $\mathcal{E}_{L}$, $\mathcal{E}_{P}$ ($\mathcal{F}$ and $\mathcal{L}$)
+17:  Obtain $f_{t}\in\mathbb{R}^{\text{num of modalities}\times\text{embedding dim}}$ using FiLM to fuse language embeddings with image and state embeddings
+18:  $F_{q}\leftarrow F_{q}\cup\{f_{t}\}$
+19:  Compute latent embedding $g_{t}=\mathcal{D}_{T}(f_{t-T},\cdots,f_{t-1},f_{t})$
+20:  Predict action $a_{t}=\mathcal{D}_{H}(g_{t})$
+21:  Interact with the environment using action $a_{t}$ and obtain done flag $d_{t+1}$
+22:  if $d_{t+1}\text{ is True}$ then break
+23:end for
+We utilize the implementations of SeqFT, ER, EWC, and PackNet provided in LIBERO (Liu et al., 2024a) and we directly use the official code of LOTUS (Wan et al., 2024).
 When training with ER and EWC, we use two GPUs, each with a batch size of 16, ensuring the total batch size remains consistent with other methods.
-For additional details on these baselines, we refer readers to [43].
-We directly use the official code of LOTUS [92].
-Since LOTUS employs a distinct architecture, we adjust its parameters to be comparable to our policy and follow the same training pipeline.
-Additionally, LOTUS utilizes the DINO-v2 [14] foundation model instead of CLIP, as recommended by the ablation study in their paper.
+Due to the unavailability of publicly released code for TAIL, we implemented it based on the description in the paper (Liu et al., 2024b).
+We integrated the code for L2M and IsCiL into our codebase, as the original implementations operate with different policies in different benchmarks.
 For LoRA-based methods, we insert LoRA modules with rank $r=8$ into the 6-th to 11-th layers of the CLIP image and text encoders, and LoRA modules with $r=16$ into the temporal transformer.
 LoRA is applied to the query and value projection matrices.
-In TAIL and IsCiL, we learn task-specific fusion modules, state encoders, and action heads, as described in [40].
-In contrast, these components are frozen in L2M following the original paper [41].
-Due to the unavailability of publicly released code for TAIL, we implemented it based on the description in the paper [40].
-We integrated the code for L2M and IsCiL into our codebase, as the original implementations operate with different policies across different benchmarks.
-For L2M, we use a pool size of 30 and the similarity loss coefficient $\lambda=0.5$.
-For IsCiL, we set the skill prototype $\mathcal{X}_{z}$ in $\mathcal{X}$ to be composed of 20 bases as in [42].
-For DMPEL, we adopt the same LoRA settings in the image encoder, text encoder, and temporal transformer, and use LoRA experts with $r=16$ for all linear layers in other submodules.
-Besides, in the Long suite, we tune both the low-rank weights $\boldsymbol{A},\boldsymbol{B}$ and biases $\boldsymbol{b}$, whereas in the Goal, Spatial, and Object suites, we only tune the low-rank weights $\boldsymbol{A},\boldsymbol{B}$.
-However, both settings lead to fewer learnable parameters than TAIL [40], which learns these sub-modules separately for each task.
+In TAIL and IsCiL, we learn task-specific fusion modules, state encoders, and action heads, as described in (Liu et al., 2024b).
+In contrast, these components are frozen in L2M following the original paper (Schmied et al., 2023).
+DMPEL (Ours).
+We adopt the same settings in the image encoder, text encoder, and temporal transformer as other LoRA-based methods, and use LoRA experts with $r=16$ for all linear layers in other sub-modules.
+Besides, in the Long suite, we tune both the low-rank weights $\bm{A},\bm{B}$ and biases $\bm{b}$, whereas in the Goal, Spatial, and Object suites, we only tune the low-rank weights $\bm{A},\bm{B}$.
+However, both settings lead to fewer learnable parameters than TAIL (Liu et al., 2024b), which learns these sub-modules separately for each task.
 We also employ dropout with probability $p=0.15$ on the coefficient for low-rank experts to avoid overfitting.
-Table 1: Environment and Policy Architecture Configuration
-Hyperparameter Value Hyperparameter Value
-Vision & Text Encoder
-Pretrained model CLIP ViT-B/16 [15] Image resolution 128*128
-Proprioceptive States
-Joint state dim. 7 Gripper state dim. 2
-Modality Fusion
-Number of hidden layers 1 Hidden size 256
-Activation GELU
-Temporal Transformer
-Number of layers 6 Number of attn. heads 8
-Embedding size 768 MLP hidden size 1024
-Max sequence Length 10 Dropout 0.15
-Action Head
-Number of hidden layers 2 Hidden size 256
-Number of Modes 5 Min. std 1e-4
-Activation Softplus Action dim. 7
-Expert router
-Number of hidden layers 2 Hidden Activation GELU
-Output Activation Sigmoid
+We provide a graphic overview of DMPEL in Figure 2 in the main text and the detailed pseudocode in Algorithm 1 (Training) and 2 (Inference).
+For further implementation specifics, please refer to the source code.
 A.2 Training Configuration
-We generally follow the setup in the LIBERO benchmark for training and evaluation [43].
-Detailed settings and hyperparameters are summarized in Table 2.
+We generally follow the setup in the LIBERO benchmark for training and evaluation (Liu et al., 2024a).
 For all experiments, we use either Nvidia A100 or H800 GPUs.
 We employ Distributed Data Parallel (DDP) for parallel training across 4 GPUs during pretraining, while using a single GPU for lifelong learning.
-We utilize the 16-bit floating point precision (FP16) training mode to accelerate the training process.
-To ensure reproducibility, we adopted three random seeds, 100, 200, 300, and we report the mean and standard deviation of the performance over three seeds throughout the paper.
-Table 2: Training Hyperparameters
-Hyperparameter Value Hyperparameter Value Hyperparameter Value
-Optimizer
-Training epochs 10 Batch size 32 Optimizer AdamW
-Betas (0.9,0.999) Learning rate 1e-4 Anneal strategy cos
-Weight decay 0.1 Gradient clip 100
-Image Augmentation
-Brightness 0.3 Saturation 0.3 Contrast 0.3
-Hue 0.3 Color jitter prob. 0.9 Rotation 15
-Translation 0.1 Affine prob. 0.6
-Evaluation
-Epochs per eval. 2 Eval. episodes 20
-A.3 Benchmark Details
-We present the language instructions for the tasks in four lifelong learning LIBERO task suites in Table 3.
-Although some tasks have similar descriptions, they are not identical due to differences in environment configurations, such as varying spatial layouts, objects, or goal positions.
-Each task provides 50 successful expert demonstrations for imitation learning.
-Table 3: Taks Instructions of LIBERO Task Suites
-Suite Instructions
+We employ 16-bit floating point precision (FP16) for training and inference to accelerate the process, with the exception of the router, which uses FP32.
+This selective precision technique is also discussed in (Fedus et al., 2022).
+To ensure reproducibility and statistical significance, we adopted three random seeds (100, 200, 300) and reported the mean and standard deviation of the performance over three seeds throughout the paper.
+We deviate from the original LIBERO benchmark by training each task for only 10 epochs (evaluating every 2 epochs), instead of 50 epochs (evaluating every 10 epochs), as we observed convergence typically within 10 epochs.
+Consequently, metrics including FWT and AUC that are averaged over epochs generally appear lower than those reported in the original LIBERO paper and some follow-up work.
+We also explored various hyperparameter settings for DMPEL, such as top-k selection, replay ratio coefficients, and weight initialization; these ablations are detailed in Figure 5.
+Otherwise, we use the same training hyperparameters across all methods and benchmarks, which proved effective.
+Hyperparameter details are summarized in Table 2.
+Table 3: Tasks Instructions of LIBERO Task Suites
+Benchmark Suite Task Instructions
 open the middle drawer of the cabinet
@@
 open the top drawer and put the bowl inside
-Goal put the bowl on top of the cabinet
+LIBERO-Goal put the bowl on top of the cabinet
 push the plate to the front of the stove
@@
 pick up the black bowl on the cookie box and place it on the plate
-Spatial pick up the black bowl in the top drawer of the wooden cabinet and place it on the plate
+LIBERO-Spatial pick up the black bowl in the top drawer of the wooden cabinet and place it on the plate
 pick up the black bowl on the ramekin and place it on the plate
@@
 pick up the bbq sauce and place it in the basket
-Object pick up the ketchup and place it in the basket
+LIBERO-Object pick up the ketchup and place it in the basket
 pick up the tomato sauce and place it in the basket
@@
 put the black bowl in the bottom drawer of the cabinet and close it
-Long put the white mug on the left plate and put the yellow and white mug on the right plate
+LIBERO-Long put the white mug on the left plate and put the yellow and white mug on the right plate
 pick up the book and place it in the back compartment of the caddy
@@
 put the yellow and white mug in the microwave and close it
+A.3 Benchmark Details
+We present the specific language instructions for the tasks in four lifelong learning LIBERO task suites in Table 3.
+Although some tasks have similar descriptions, they are not identical due to differences in environment configurations, such as varying spatial layouts, objects, or goal positions.
+Each task provides 50 successful expert demonstrations for imitation learning.
 Appendix B Additional Results
-B.1 Learning curves
-We present the learning curves for all methods in four lifelong learning LIBERO task suites in Figure 7-10 .
+B.1 Ablation Studies
+((a))
+((b))
+((c))
+Figure 8: Ablation studies in the LIBERO-Goal suite. (a) Different task orders (b) Different rank sizes (c) Different number of demonstrations.
+Figure 9: Ablation studies on router design in the LIBERO-Goal suite.
+We further investigate the performance of DMPEL across various settings.
+First, we assess its robustness to different task orders, as shown in Figure 8(a).
+The default order provided by the LIBERO benchmark is 0123456789, the reverse order is 9876543210, and a randomly generated task order is 4687312095.
+Additionally, we analyze the impact of the rank size of each LoRA expert and the number of demonstrations for each new task, illustrated in Figures 8(b) and 8(c), respectively.
+We utilize the Sigmoid function to constrain the router output instead of applying Softmax normalization.
+As noted in the related work, DMPEL is heavily inspired by research on model fusion and merging, such as SD-LoRA (Wu et al., 2025) and LoRAHub (Huang et al., 2024), which demonstrate that adjusting only the magnitude of LoRA can yield promising transfer results.
+The interval [0, 2] serves as a hyperparameter based on intuition to limit the magnitude, with 1 as the midpoint, representing the direct use of the previous LoRA module without magnitude alteration.
+We empirically compare the performance of Sigmoid and Softmax with various top-k strategies, as shown in Figure 9.
+Notably, router with Sigmoid generally outperforms the ones with Softmax, and the top-1 with Softmax variant exhibited the lowest performance.
+We hypothesize that using raw coefficients to modulate the magnitudes of LoRA experts provides greater flexibility, as these coefficients are independent for each expert.
+In contrast, Softmax normalization enforces that the sum of coefficients equals 1, meaning that increasing one coefficient necessarily decreases another.
+During expert coefficient replay, we apply the MSE loss between old and new coefficients in all cases for its simplicity.
+When transitioning from task $k$ to task $k+1$, we pad the coefficient vector with 0 to maintain alignment with the current library size and ensure that later introduced experts are excluded from previous tasks.
+We hypothesize that minimizing the discrepancy of logits may be a more effective approach for the top-1 Softmax variant due to its one-hot property.
+In this context, padding zeros to the coefficients becomes padding -$\infty$ to the logits.
+We leave further investigation of this for future work.
+Beyond the experiment on the standard $K=10$ task sequence form the LIBERO benchmark, we also investigate the applicability of DMPEL to ultra-long sequences.
+To this end, we concatenate the LIBERO-Object, Goal, and Spatial benchmark to form a task sequence with $K=30$.
+As mentioned in Section 5.2 and illustrated in Figure 6(b), we prune underactivated low-rank experts every 10 tasks to control the growth of the library.
+The results shown in Figure 10 indicate that DMPEL consistently outperforms FFT with experience replay in success rate, while incurring significantly lower computational and storage costs (see Figure 1).
+However, we observe a gradual decline in performance as the number of tasks increases.
+Given the small size of the router, this is not surprising.
+When the number of distinct tasks continue to increase, the router may be exhausted.
+One potential solution is to progressively scale up the router as the task count increases.
+Since all expert coefficients are stored in the replay buffer, we should be able to recover the coefficient distribution in the enlarged router.
+We plan to explore this in future research.
+Figure 10: Average success rate on encountered tasks in the ultra long task sequence concatenated with LIBERO-Object, Goal, and Spatial benchmarks.
+B.2 Visualization Analysis
+We plot some representative trajectories from different benchmark suites in Figure 11-13.
+In Figure 11, we observe that the temporal decoder reuses the previously learned low-rank experts throughout the entire trajectory, while the state encoder relies on a newly introduced expert, especially when grasping the bowl.
+In Figure 12, we can see the switching process between different experts as the robot approaches, grasps, transports, and places down the bowl.
+Both the vision and state encoder relies on the new expert.
+In Figure 13, the vision encoder reuses previously learned expert but the fusion sub-module learns a new expert to effectively fuse features for manipulation.
+((a))
+((b))
+Figure 11: Visualization Analysis on task 9 from LIBERO-Goal task: put the bowl on the plate.
+((a))
+((b))
+Figure 12: Visualization Analysis on task 7 from LIBERO-Spatial task: pick up the black bowl next to the cookie box and place it on the plate.
+((a))
+((b))
+Figure 13: Visualization Analysis on task 5 from LIBERO-Object task: pick up the ketchup and place it in the basket.
+B.3 Learning curves
+We present the learning curves for all methods in four lifelong learning LIBERO task suites in Figure 14-17 .
 The $y$-axis indicates the success rate, while the $x$-axis depicts the agent’s training process over the course of lifelong learning.
-For example, the Task 1 subplot in the figure shows the agent’s performance on the first task as it learns ten tasks sequentially.
+For example, the Task 1 subplot in the figure shows the agent’s performance on the first task when it learns ten tasks sequentially.
+Learning curves demonstrate that DMPEL achieves superior forward transfer with reduced forgetting compared to baselines.
 Figure 14: Visualization of learning curves on the LIBERO-Goal benchmark.
@@
 Figure 17: Visualization of learning curves on the LIBERO-Long benchmark.
+Appendix C The Usage of Large Language Models
+We only use Large Language Models (LLMs) to assist and polish writing based on prompts such as "Please help me polish the paragraph and check for any grammatical mistakes."